### v1

# Fast nonparametric feature selection with error control using integrated path stability selection

Abstract.
Feature selection can greatly improve performance and interpretability in machine learning problems. However, existing nonparametric feature selection methods either lack theoretical error control or fail to accurately control errors in practice. Many methods are also slow, especially in high dimensions. In this paper, we introduce a general feature selection method that applies integrated path stability selection to thresholding to control false positives and the false discovery rate. The method also estimates $q$-values, which are better suited to high-dimensional data than $p$-values. We focus on two special cases of the general method based on gradient boosting (IPSSGB) and random forests (IPSSRF). Extensive simulations with RNA sequencing data show that IPSSGB and IPSSRF have better error control, detect more true positives, and are faster than existing methods. We also use both methods to detect microRNAs and genes related to ovarian cancer, finding that they make better predictions with fewer features than other methods.
1.  Introduction
Identifying the important features in a dataset can greatly improve performance and interpretability in machine learning problems (Theng and Bhoyar, 2024). For example, in genomics, often only a small fraction of genes (features) are related to a disease of interest (response). By identifying these genes, scientists can save time and resources while gaining insights that would be difficult to uncover otherwise (Theng and Bhoyar, 2024). A common form of feature selection is thresholding: Given a function $\Phi$ that assigns a degree of importance $\Phi(j)$ to each feature $j$ in a dataset, one seeks a value $\lambda$ such that the selected set $\hat{S}_{\lambda}=\{j:\Phi(j)\geq\lambda\}$ contains important features and its complement, $\hat{S}^{c}_{\lambda}=\{j:\Phi(j)<\lambda\}$, does not. Choosing $\lambda$ can be challenging, and many methods for doing so come with few theoretical guarantees (Huynh-Thu et al., 2012).
Stability selection is a popular technique for improving the performance of feature selection algorithms (Meinshausen and Bühlmann, 2010). However, its implementation relies on relatively weak upper bounds on the expected number of false positives, E(FP), that is, on the expected number of unimportant features that are selected. Recently, Melikechi and Miller (2024) proved that much stronger bounds hold for integrated path stability selection (IPSS), which consequently identifies more true positives in practice than stability selection. Until now, IPSS has only been applied to parametric linear models, and its generalization to more complex models has not been explored.
In this work, we extend IPSS to nonlinear, nonparametric models by applying it to thresholding. The result is a feature selection method with tight upper bounds on E(FP) that are characterized by novel quantities called efp scores. In addition to controlling E(FP), efp scores also approximately control the false discovery rate (FDR) and estimate the $q$-values for each feature, which are more reliable and robust than $p$-values in genomics and other high-dimensional settings (Storey and Tibshirani, 2003). Lastly, by integrating over threshold values, IPSS for thresholding replaces the problem of specifying a single threshold $\lambda$ with the simpler task of specifying either a target E(FP) or a target FDR.
We develop two specific instances of the general method: IPSS for gradient boosting (IPSSGB) and IPSS for random forests (IPSSRF). In simulations, we find that IPSSGB has better false positive control and identifies more true positives than nine other feature selection methods, and that both IPSSGB and IPSSRF outperform IPSS for linear models when the linearity assumptions of the latter are violated. More generally, IPSSGB and IPSSRF control E(FP) and FDR more accurately than other stability selection and knockoff-based methods with theoretical error control. In Section 5, we find that IPSSGB and IPSSRF effectively identify microRNAs and genes related to ovarian cancer, achieving better predictive performance than other feature selection methods while using fewer features. Both also run faster than leading methods.
Organization. In Section 2, we introduce efp scores, IPSS for thresholding in general, and IPSS for gradient boosting and random forests. In Section 3, we describe several leading feature selection methods, which are compared to IPSSGB and IPSSRF in a simulation study in Section 4. In Section 5, we analyze ovarian cancer data. We conclude in Section 6 with a discussion.
2.  Methods
We introduce efp scores (Section 2.1), review IPSS (Section 2.2), introduce IPSS for thresholding (Section 2.3), and apply IPSS to gradient boosting and random forests (Section 2.4).
[…]
2.1.  efp scores.
Suppose $S\subseteq\{1,\ldots,p\}$ is an unknown subset of true important features that we wish to estimate using $Z_{1:n}$. An efp (expected false positive) score is a function $\mathtt{efp}_{Z_{1:n}}:\{1,\dots,p\}\to[0,\infty)$ that depends on $Z_{1:n}$ and satisfies the following:
For all $t\geq 0$, if $\hat{S}(t)=\{j:\mathtt{efp}_{Z_{1:n}}(j)\leq t\}$ then $\mathrm{E}(\mathrm{FP}(t))\leq t$,
where $\mathrm{E}(\mathrm{FP}(t))=\mathrm{E}\lvert\hat{S}(t)\cap S^{c}\rvert$ is the expected number of false positives in $\hat{S}(t)$. That is, the estimator $\hat{S}(t)$ of $S$ selects at most $t$ false positives on average. A trivial example of an efp score is $\mathtt{efp}_{Z_{1:n}}(j)=p$ for all $j$. This corresponds to selecting either no features or all features. Specifically, if $t\in[0,p)$, then $\hat{S}(t)=\varnothing$ and $\mathrm{E}(\mathrm{FP}(t))=0\leq t$, while if $t\in[p,\infty)$, then $\hat{S}(t)=\{1,\dots,p\}$ and $\mathrm{E}(\mathrm{FP}(t))\leq t$, since the number of false positives is at most $p$.
The quality of an efp score is measured by the tightness of its bounds $\mathrm{E}(\mathrm{FP}(t))\leq t$. Better efp scores have tighter bounds because tight bounds enable accurate false positive control via the parameter $t$. Accurate control in turn leads to more true positives in $\hat{S}(t)$ since weak bounds overestimate the number of false positives, reducing the total number of features selected.
E(FP) and efp scores are related to two other quantities of significant interest: the false discovery rate (FDR) and $q$-values. Informally, the false discovery rate is the expected ratio between the number of false positives and the total number of features selected, $\mathrm{FDR}=\mathrm{E(FP/(TP+FP))}$, and the $q$-value of feature $j$ is the smallest FDR when $j$ is selected (Storey, 2003). When $p$ is large, as is often the case in genomics, Storey (2003) showed that
$\begin{split}\mathrm{pFDR}(t)&\approx\mathrm{FDR}(t)\approx\frac{\mathrm{E}(\mathrm{FP}(t))}{\mathrm{E}\lvert\hat{S}(t)\rvert}\leq\frac{t}{\mathrm{E}\lvert\hat{S}(t)\rvert},\end{split}$ (2.1)
where $\mathrm{pFDR}(t)=\mathrm{E}(\mathrm{FP}(t)/\lvert\hat{S}(t)\rvert\mid\lvert\hat{S}(t)\rvert>0)$ is the positive false discovery rate, and the inequality holds by the definition of an efp score. It follows that the $q$-value of feature $j$ satisfies
$\begin{split}q_{j}&=\inf_{\{t:\mathtt{efp}(j)\leq t\}}\mathrm{pFDR}(t)\lesssim\inf_{\{t:\mathtt{efp}(j)\leq t\}}\frac{t}{\mathrm{E}\lvert\hat{S}(t)\rvert},\end{split}$ (2.2)
where the equality is the definition of the $q$-value (here, $\mathtt{efp}(j)$ denotes the observed value of the test statistic $\mathtt{efp}_{Z_{1:n}}(j)$) (Storey, 2003), and the approximate inequality $\lesssim$ holds by Equation 2.1. Thus, when the efp score has tight bounds, the $q$-value of feature $j$ is well-approximated by the rightmost term in Equation 2.2, which is easily estimated in practice by replacing $\mathrm{E}\lvert\hat{S}(t)\rvert$ with $\lvert\hat{S}(t)\rvert$. Similarly, by Equation 2.1, $\mathrm{FDR}(t)$ is approximately bounded by $t/\lvert\hat{S}(t)\rvert$. So, as an alternative to specifying the target E(FP) parameter $t$, one can control the FDR at level $\alpha$ by choosing the largest set $\hat{S}(t)$ such that $t/\lvert\hat{S}(t)\rvert\leq\alpha$. The largest such set is chosen to maximize true positives.
2.2.  Integrated path stability selection.
In Section 2.1, we showed that efp scores can be used to control E(FP), the FDR, and $q$-values. In this section, we review integrated path stability selection (IPSS), which constructs efp scores from arbitrary feature selection algorithms. The advantage of IPSS over other stability selection methods is that its efp scores have the tightest E(FP) bounds (Melikechi and Miller, 2024).
Let $S$ be the unknown subset of important features as before, and let $\hat{S}_{\lambda}\subseteq\{1,\ldots,p\}$ be an estimator of $S$ that depends on the data and a parameter $\lambda>0$, with larger values of $\lambda$ corresponding to fewer features being selected. Importantly, $\hat{S}_{\lambda}$ and $\hat{S}(t)$ are distinct estimators of $S$: The former is a baseline selection algorithm whose parameter $\lambda$ will always appear as a subscript, and the latter is an estimator based on efp scores whose parameter $t$ will always appear as a functional argument.
Stability selection uses subsampling to construct $\hat{S}(t)$ from $\hat{S}_{\lambda}$ (Meinshausen and Bühlmann, 2010; Shah and Samworth, 2013). Specifically, rather than estimate $S$ using all of the data at once, one randomly draws disjoint subsets $A_{1},A_{2}\subseteq\{1,\dots,n\}$ of size $\lfloor n/2\rfloor$ and evaluates $\hat{S}_{\lambda}(Z_{A_{1}})$ and $\hat{S}_{\lambda}(Z_{A_{2}})$ at all $\lambda$ in some interval $\Lambda\subseteq(0,\infty)$, where $Z_{A}=(Z_{i}:i\in A)$. After $B$ subsampling iterations, the estimated selection probability $\hat{\pi}_{j}(\lambda)$ is the proportion of times feature $j$ is selected over all $2B$ subsets, $\hat{\pi}_{j}(\lambda)=\frac{1}{2B}\sum_{b=1}^{2B}\operatorname{\mathds{1}}(j\in\hat{S}_{\lambda}(Z_{A_{b}}))$.
Melikechi and Miller (2024) prove that for any $\Lambda\subseteq(0,\infty)$, any probability measure $\mu$ on $\Lambda$, and certain functions $f:[0,1]\to\mathbb{R}$, the function $\mathtt{efp}_{Z_{1:n}}:\{1,\dots,p\}\to[0,p]$ defined by
$\mathtt{efp}_{Z_{1:n}}(j)$ $=\min\left\{\frac{\mathcal{I}(\Lambda)}{{\int}_{\Lambda}f(\hat{\pi}_{j}(\lambda))\mu(d\lambda)},\ p\right\},$ (2.3)
is a valid efp score, where $\mathcal{I}(\Lambda)$ is a parameter that is described below and in Section S1. While several choices of $f$ are possible in Equation 2.3, we always use $f(x)=(2x-1)^{3}\operatorname{\mathds{1}}(x\geq 0.5)$ because the resulting efp score bounds $\mathrm{E}(\mathrm{FP}(t))\leq t$ are the tightest of any existing version of stability selection (Melikechi and Miller, 2024).
[…]
2.3.  IPSS for thresholding.
An importance function is a (possibly random) map $\Phi_{Z_{1:n}}:\{1,\dots,p\}\to\mathbb{R}$ that uses the data to assign an importance score $\Phi_{Z_{1:n}}(j)$ to each feature. The possible randomness, which is additional to the randomness in $Z_{1:n}$, can come from, for example, random subsampling in tree-based algorithms. Suppressing $Z_{1:n}$ from the notation for now, assume that $\Phi(j)<\Phi(k)$ means $j$ is less important than $k$ according to $\Phi$. For example, in linear regression where $Z_{i}=(X_{i},Y_{i})$ with $Y_{i}=\beta^{\mathtt{T}}X_{i}+\epsilon_{i}$ and $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$, a viable importance function is the magnitude of the estimated regression coefficient, $\Phi(j)=\lvert\hat{\beta}_{j}\rvert$. Another class of importance functions—the focus in this work—come from tree ensemble methods such as gradient boosted trees and random forests, where again we are in the supervised setting $Z_{i}=(X_{i},Y_{i})$ (Friedman, 2001; Breiman, 2001). Given a collection of binary decision trees, $\mathcal{T}$, these importance functions take the form
$\Phi(j)$ $=\frac{1}{\lvert\mathcal{T}\rvert}\sum_{T\in\mathcal{T}}\sum_{v\in T}\Delta\varphi(v)\operatorname{\mathds{1}}(j=j_{v})$ (2.4)
where the outer sum is over all trees $T\in\mathcal{T}$, the inner sum is over all nodes $v\in T$, $j_{v}$ is the feature used to split node $v$, and $\varphi(v)$ measures the impurity of $v$. For regression, we use the squared error loss, $\varphi(v)=\sum_{i\in v}(Y_{i}-\overline{Y}_{v})^{2}/\lvert v\rvert$, where $v\subseteq\{1,\dots,n\}$ is identified with the subset of samples in node $v$, and $\overline{Y}_{v}=\sum_{i\in v}Y_{i}/\lvert v\rvert$ is the empirical mean of the responses in $v$. For binary classification, we use the Gini index, $\varphi(v)=2p_{0}(v)p_{1}(v)$ where, for $a\in\{0,1\}$, $p_{a}(v)=\sum_{i\in v}\operatorname{\mathds{1}}(Y_{i}=a)/\lvert v\rvert$ is the proportion of responses in $v$ that equal $a$. The change in impurity
$\Delta\varphi(v)$ $=\varphi(v)-\bigg{(}\frac{\lvert v_{L}\rvert}{\lvert v\rvert}\varphi(v_{L})+\frac{\lvert v_{R}\rvert}{\lvert v\rvert}\varphi(v_{R})\bigg{)}$
is the impurity difference between $v$ and its children, $v_{L}$ and $v_{R}$. Large positive values of $\Delta\varphi(v)$ indicate that the feature $j_{v}$ used to split node $v$ successfully partitions the data in a manner that is consistent with the objective of the tree. Importance scores naturally lend themselves to thresholding: Given a $\Phi$ and a threshold $\lambda$, the selected set of features is $\hat{S}_{\lambda}=\{j:\Phi(j)\geq\lambda\}$.
Algorithm S1 outlines IPSS for thresholding. The main steps are as follows: First, features are preselected according to the procedure described in Section S1.4. This is a common preliminary step in many feature selection algorithms. Next, importance scores are evaluated for the preselected features using random, disjoint halves of the data. This process is repeated $B$ times, yielding $2B$ importance scores for each feature. We then construct a grid of $\lambda$ values. The largest, $\lambda_{\mathrm{max}}$, is the maximum importance score across all features and all $2B$ sets of scores (hence, $\hat{S}_{\lambda_{\mathrm{max}}}=\varnothing$). Next, starting from $\lambda_{\mathrm{min}}=\lambda_{\mathrm{max}}$, we decrease $\lambda_{\mathrm{min}}$ one step at a time, usually on a log scale, iteratively updating the integral $\mathcal{I}([\lambda_{\mathrm{min}},\lambda_{\mathrm{max}}])$ at each step in the form of a Riemann sum approximation until $\mathcal{I}([\lambda_{\mathrm{min}},\lambda_{\mathrm{max}}])$ surpasses the cutoff $C=0.05$. Once $C$ is surpassed, the $\mathtt{while}$ loop stops and the feature-specific selection probabilities and integrals are evaluated. The algorithm outputs the efp scores for each feature. Given a target E(FP), $t$, the set of selected features is $\hat{S}(t)=\{j:\mathtt{efp}_{Z_{1:n}}(j)\leq t\}$.
2.4.  IPSSGB and IPSSRF.
We consider two instances of the general method: IPSSGB uses gradient boosting for the baseline algorithm $\hat{S}_{\lambda}$, and IPSSRF uses random forests for $\hat{S}_{\lambda}$. All other IPSS parameters are identical, except we use $B=50$ for IPSSRF and $B=100$ for IPSSGB. The latter led to slightly more stable results with hardly any increase in runtimes (no such improvement was observed for IPSSRF).
In all that follows, random forests are implemented with scikit-learn (Pedregosa et al., 2011) and gradient boosting with XGBoost (Chen and Guestrin, 2016). For random forests, all parameters are set to their default settings except for the proportion of features considered when splitting each node (called $\mathtt{max\_features}$ in scikit-learn and often $\mathtt{mtry}$ elsewhere), which we change from $1$ to $1/3$. This is a standard choice (Biau and Scornet, 2016) and led to an approximately $3$-fold speedup in IPSSRF without any discernible effect on the results. For gradient boosting, we also change the proportion of features considered at each split from $1$ to $1/3$. Aside from this, the only non-default parameter is the maximum depth of each tree, which we change from $6$ to $1$, making each tree a stump. This significantly improved the performance of IPSSGB, both in terms of speed and feature selection. Finally, while many different importance functions exist for gradient boosting and random forests, we use mean decrease impurity, Equation 2.4, with the squared error loss for regression and the Gini index for classification, as defined in Section 2.3. Experimentation with other importance functions showed similar results, though a more detailed study of IPSS with different importance functions is a potentially direction for future work.
2.4.1.  Further background.
The seminal work on gradient boosting is Friedman (2001). For random forests, we refer to the original work of Breiman (2001), and to Biau and Scornet (2016) for an excellent review with a detailed description of importance scores. Gradient boosting, random forests, and their importance scores are also detailed in Hastie et al. (2009). Finally, our formulation of tree-based scores partly follows Louppe et al. (2013), who provide a clear and concise formulation of these quantities.
3.  Other methods
We describe several feature selection methods that are compared to IPSSGB and IPSSRF in Sections 4 and 5. We selected these methods due to their popularity and superior performance relative to other methods in several comparison studies (Degenhardt et al., 2019; Sanchez-Pinto et al., 2018; Speiser et al., 2019).
3.1.  Methods with error control.
The closest method to IPSSGB in terms of its underlying approach is that of Hofner et al. (2015), referred to here as SSBoost. Unlike IPSSGB—which uses importance scores from gradient boosting—SSBoost applies stability selection to choose the number of features used per boosting run. Furthermore, IPSSGB uses IPSS to construct efp scores, whereas SSBoost uses a version of stability selection introduced by Shah and Samworth (2013). This is perhaps the most significant difference since the efp scores for IPSS have much tighter bounds than those for other forms of stability selection (Melikechi and Miller, 2024). From a practical standpoint, this causes other versions of stability selection to identify fewer important features than IPSS for the same target E(FP).
IPSS has previously been used by Melikechi and Miller (2024) with lasso for regression and $\ell^{1}$-regularized logistic regression for classification (Tibshirani, 1996; Lee et al., 2006). These methods, abbreviated as IPSSL, are part of a more general class of $\ell^{1}$-regularized estimation techniques to which stability selection is commonly applied (Meinshausen and Bühlmann, 2010), namely,
$\hat{\beta}$ $=\operatorname*{arg\,min}_{\beta\in\mathbb{R}^{p}}\sum_{i=1}^{n}\mathcal{L}(Y_{i},\beta^{\mathtt{T}}X_{i})+\lambda\sum_{j=1}^{p}\lvert\beta_{j}\rvert,$
where $\mathcal{L}$ is a log-likelihood function and $\lambda>0$ controls the strength of the $\ell^{1}$ penalty. By contrast, since decision trees are nonparametric, neither IPSSGB nor IPSSRF assume an underlying model. Furthermore, unlike linear regression and logistic regression, decision trees can effectively capture nonlinear relationships in data. The nonlinear and nonparametric nature of decision trees motivated our extension of IPSS to gradient boosting and random forests. In Section 4, we find that both methods significantly outperform IPSSL in the presence of nonlinearity.
Another method to which we compare is knockoff boosted trees (KOBT) (Jiang et al., 2021). KOBT applies the model-X knockoff framework of Candes et al. (2018) to boosted trees to obtain a method with theoretical FDR control. We use the R package KOBT, preselecting $\mathtt{bound}=200$ features and generating $\mathtt{num}=100$ knockoff matrices. The knockoff type is shrunken Gaussian. These choices were made to balance performance and runtime. Computation times were much longer for larger values of $\mathtt{bound}$ or $\mathtt{num}$, or when using a different version of KOBT that employs principal components to avoid the Gaussian assumption. Moreover, we found that these computationally more expensive choices did not noticeably improve performance or FDR control.
We also tested the method of Coleman et al. (2022), which achieves theoretical error control by using random forests for hypothesis testing. However, one test run with default parameters on simulated data with 500 samples, 500 features, and 20 true features took over 51 minutes, returning 15 true positives and 43 false positives. By contrast, IPSSGB with a target E(FP) of 2 took 11 seconds and returned 10 true positives and 0 false positives on the same data. This method is omitted from our study because its performance does not appear to justify its excessive runtime.
3.2.  Methods without error control.
We also compare to several popular methods without theoretical error control: Boruta (Kursa et al., 2010), random forests (RF) (Breiman, 2001), Vita (Janitza et al., 2018), VSURF (Genuer et al., 2010), and XGBoost (Chen and Guestrin, 2016). Boruta, Vita, and VSURF use random forests. In terms of performance, Speiser et al. (2019) found Boruta and VSURF were among the best out of $13$ random forest-based methods, and Degenhardt et al. (2019) found Boruta and Vita outperformed $5$ other methods on several metrics.
Method Package Error control Base method Non-default settings
IPSSGB ipss (Python) Boosting
IPSSRF ipss (Python) Random forest
IPSSL ipss (Python) Lasso
SSBoost XGBoost (Python) Boosting $\tau=0.75$
with stabs (R)
KOBT KOBT (R) Boosting $\texttt{num}=100$, $\texttt{bound}=200$,
[…]
Boruta Boruta (R) Random forest
RF scikit-learn (Python) Random forest
Vita vita (R) Random forest $p$-value threshold $=0$
VSURF VSURF (R) Random forest VSURF_pred
XGBoost XGBoost (Python) Boosting
Table 1. Feature selection methods. Packages, when available, are listed with the language used to implement them in parentheses. Non-default settings are elaborated upon in Sections 3.1 and 3.2. For methods with no non-default settings, we use the default settings in their respective packages.
Table 1 summarizes the methods. Details regarding the implementation and choice of settings are as follows. Hofner et al. (2015) provide code for SSBoost that combines the R packages mboost (Hofner et al., 2014) and stabs (Hofner and Hothorn, 2017) in the form of a worked example, but the mboost implementation of boosting was prohibitively slow in the dimensions we consider. By adapting their code to use XGBoost in place of mboost for boosting and by preselecting features exactly as we do for IPSSGB (Section S1.4), we were able to reduce SSBoost runtimes considerably with no apparent change in results. The XGBoost parameters used for SSBoost are the same as those used for IPSSGB; specifically, all XGBoost parameters are set to their default values except the proportion of features considered at each split (changed from 1 to 1/3) and the depth of each tree (changed from 6 to 1). For the stability selection part of SSBoost we use the default parameters in stabs, and the selection threshold is set to $\tau=0.75$, which is the middle of the interval $(0.6,0.9)$ recommended by Meinshausen and Bühlmann (2010). For Vita, we set the $p$-value threshold to 00, and for VSURF we use the function VSURF_pred rather than VSRUF_interp to select the final set of features (Genuer et al., 2010); both choices correspond to as few features as possible being selected under the constraints of the respective algorithms, a favorable quality due to the sparsity in our forthcoming studies. For XGBoost and RF, we run the respective algorithms on the full dataset with default parameters, then select the top ranked features by their importance scores.
4.  Simulation study
In this section, we conduct a simulation study to evaluate the performance of the methods in Table 1 when the true set of important features is known. To make the simulations more realistic, we use features from real data rather than generating them from known distributions. Specifically, given $n$ and $p$, in each trial we randomly select $n$ samples and $p$ features from RNA sequencing (RNA-seq) data from 569 ovarian cancer patients (samples) and 6426 genes (features) (Vasaikar et al., 2018). This publicly available dataset, part of The Cancer Genome Atlas (Weinstein et al., 2013), was chosen because it is high dimensional and the features follow a variety of distributions. For instance, Figure 2 shows the standardized empirical distributions of five randomly selected genes, which, from left to right, are relatively flat, skewed left, approximately Gaussian, skewed right, and contain outliers. Furthermore, the genes exhibit complex correlation structures, with maximum and average absolute pairwise correlations of approximately $0.95$ and $0.17$ after standardization, respectively.
[…]
Figure 1. Simulating the response. Random rows and columns are selected from the full RNA-seq dataset to create an $n$-by-$p$ design matrix $X$, whose columns are then split into important features, $X_{S}$, and unimportant features, $X_{S^{c}}$. The columns of $X_{S}$ are further partitioned into $G$ groups; the above figure shows the case $G=3$. The features in each group are summed to obtain $\xi_{g}$, and a different realization $f_{\theta_{g}}$ of $f_{\theta}$ is applied to $\xi_{g}$ for each $g$. The response is the sum of the resulting group-specific signals, plus noise.
Figure 2. Distributions of five genes from the RNA-seq dataset.
4.1.  Simulation details.
Algorithm S2 describes the simulation procedure, which is also illustrated in Figure 1. In all steps, “randomly select” means select a parameter uniformly at random from its domain. The general outline is as follows: First, randomly select an $n$-by-$p$ submatrix $X$ of the full RNA-seq dataset and standardize its columns to have mean 0 and variance 1. Next, randomly select a true subset of $p_{\mathrm{true}}$ important features $S$, and partition it into $G$ groups, $S_{1},\dots,S_{G}$. A different realization of a randomized nonlinear function $f_{\theta_{g}}$, defined in Equation S2.1, is applied to the standardized sum $\xi_{g}=(\hat{\xi}_{g}-\bar{\xi}_{g})/\bar{\sigma}_{g}$ of the features in each group $S_{g}$, where $\bar{\xi}_{g}$ and $\bar{\sigma}_{g}$ are the empirical mean and standard deviation of $\hat{\xi}_{g}=\sum_{j\in S_{g}}X_{j}$. The resulting values are summed over all groups to generate a signal $\eta=\sum_{g=1}^{G}f_{\theta_{g}}(\xi_{g})$, and noise is added to this signal to generate a response $y$. This scheme produces data with highly complex interactions between features and the response, going well beyond the additive setting $y=\sum_{j\in S}f_{j}(X_{j})+\epsilon$.
Regression and classification data are simulated with $p=500$, $2000$, and $5000$ features for a total of six experiments. Each experiment consists of $100$ trials, outlined in Algorithm S2. Each trial uses $n=500$ samples and $\lvert S\rvert=p_{\mathrm{true}}$ important features. For regression, the response is $y=\sum_{g=1}^{G}f_{\theta}(\xi_{g})+\epsilon$, where $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$ and the variance $\sigma^{2}$ is selected according to a specified signal-to-noise ratio (SNR). For classification, we draw $y_{i}\sim\mathrm{Bernoulli}(\pi_{i})$ where $\pi_{i}=1/(1+\exp(-u\eta_{i}))$. A new $p_{\mathrm{true}}$ and SNR (or $u$), as well as new sets of samples, features, and important features, are drawn prior to each trial, with $p_{\mathrm{true}}$, SNR, and $u$ chosen according to Table S1. The many sources of randomness in our setup are introduced so that the simulations cover a wide range of settings.
4.2.  Simulation results.
Figures 3, LABEL:, 4, LABEL:, and 5 show the results. The empirical false discovery and true positive rates are $\mathrm{FDR}=\mathrm{FP}/(\mathrm{TP}+\mathrm{FP})$ and $\mathrm{TPR}=\mathrm{TP}/(\mathrm{TP}+\mathrm{FN})$, where FP, TP, and FN are the number of false positives, true positives, and false negatives, respectively. The upper left (high TP or TPR and low FP or FDR) corresponds to better performance in Figures 3 and 4. The ability of the different methods to control E(FP) (Figure 3) and FDR (Figure 4) at their target values is shown when applicable. Specifically, SSBoost and KOBT provide theoretical control of E(FP) and FDR, respectively, while the IPSS methods can be specified to control either E(FP) or FDR (Section 2.2). In Figure 3, the dashed vertical line signifies a target E(FP) of 3, with symbols showing the average FP and TP for the relevant methods at this target value. In particular, the for a method with perfect E(FP) control should lie on the dashed vertical line. Similarly, the and symbols in Figure 4 correspond to target FDRs of 0.1 and 0.3, respectively.
IPSSGB has the best performance overall. Its curves are the farthest to the upper left in every plot in Figures 3 and 4, and it has the best control of both E(FP) and FDR. In Figure 3, for example, the for IPSSGB is the closest to the vertical dashed line in all experiments, and similarly in Figure 4. Furthermore, IPSSGB identifies the most true positives, both in terms of average TP and average TPR. Notably, at a target E(FP) of 3 and a target FDR of 0.3, IPSSGB has the largest average TP and highest TPR in nearly every setting, while having significantly fewer false positives than most other methods. Figure 5 also shows that IPSSGB is one of the fastest methods, running in under 20 seconds in all settings.
Among the other methods with theoretical error control, IPSSRF performs well in terms of identifying true positives while controlling false positives, though not as well as IPSSGB. IPSSL, whose model assumptions are violated, performs poorly, neither controlling false positives at the target E(FP) or FDR levels, nor identifying many true ones. SSBoost, whose curves are obtained by varying the target E(FP) between 0 and 10, has few false positives but also few true positives. Observe, for example, that the for SSBoost in Figure 3 is always significantly lower than those for IPSSGB and IPSSRF, despite all three methods having the same target E(FP). This is due in part to the weakness of the efp scores used by SSBoost relative to those used by IPSS (Melikechi and Miller, 2024). KOBT performs poorly in general; most notably, its actual FDRs far exceed its target FDRs of 0.1 and 0.3 in all of the simulation settings in Figure 4. KOBT also runs in around 90 seconds, making it one of the slowest methods. With the exception of IPSSL for classification, all of the stability selection-based methods run in at most 40 seconds, often less.
The absence of error control for Boruta and Vita is clearly apparent in Figures 3 and 4. Both methods usually have FDRs over 0.75—often 20 or more false positives on average—far surpassing IPSSGB and IPSSRF. Despite this, Boruta and Vita identify fewer true positives than IPSSGB with a target E(FP) of 3 and, in many cases, IPSSGB with a target FDR of 0.3. The RF and XGBoost curves are obtained by increasing the number of features selected, as ranked by their importance scores, from 0 to 30; that is, the first nonzero point along the curve is obtained by keeping only the feature with the highest importance score, the second point is obtained by keeping the features with the two largest importance scores, and so on. Interestingly, RF and XGBoost significantly underperform IPSSGB and IPSSRF in terms of true and false positives across the whole 0 to 30 range, indicating that IPSS is able to identify important features that simple thresholding of the baseline algorithm cannot, no matter how many features are kept. In terms of runtime, RF and XGBoost run quickly except in the $p=5000$ regression experiment. Boruta and Vita are typically slower than the IPSS methods, and this disparity grows with the number of features.
Figure 3. TP versus FP for the six simulation experiments. Columns correspond to the number of features, $p$, and the top and bottom rows correspond to regression and classification, respectively. Curves for the three IPSS methods and SSBoost are obtained by varying the target E(FP), with showing results when the target E(FP) is 3. The KOBT curve is obtained by varying the target FDR, and the RF and XGBoost curves are obtained by increasing the number of features selected (as ranked by their importance scores) from 0 to 30. The remaining methods, indicated by the symbols, do not have curves because they select fixed sets of features.
Figure 4. TPR versus FDR for the six simulation experiments. Results are similar to those in Figure 3, except now we show the FDR and TPR rather than E(FP) and E(TP), and curves for the three IPSS methods and KOBT are obtained by varying the target FDR, with and showing results when the target FDR is 0.1 and 0.3.
Figure 5. Runtimes for the six experiments. Columns correspond to the number of features, $p$, and the top and bottom rows correspond to regression and classification, respectively. Orange lines denote median runtime and whiskers extend to 1.5 times the interquartile range.
5.  Ovarian cancer
We apply IPSSGB, IPSSRF, and the other methods in Table 1 to microRNA (miRNA) data from the cohort of ovarian cancer patients described in Section 4 (Vasaikar et al., 2018). We report two studies here; see Section S3 for results from additional studies, including analyses of the RNA-seq data from this cohort. In the first study, the features are the 588 miRNAs and the response variable is prognosis, which is defined as whether the patient was still alive at the time of last follow-up. In the second study, the response is one of the miRNAs, miR-150, identified as important in the first stage, and the features are the remaining miRNAs. Thus, since miRNA measurements are continuous, the first study is a binary classification problem, and the second study is regression.
All analyses, both here and in Section S3, are conducted using 10-fold cross-validation (CV) as follows. First, all patients with missing values are removed, resulting in $n=442$ and $n=453$ for the first and second studies, respectively. The remaining patients are evenly partitioned into 10 groups. In each of the 10 CV steps and for each selection method, one group of patients is set aside (the test set), and a set of miRNAs is selected by applying the method to the data in the remaining groups (the training set). For the IPSS methods, we also compute the efp scores and $q$-values of each miRNA on the training set. Next, for each method we construct three predictive models—a linear model (either lasso or $\ell^{1}$-regularized logistic regression), a random forest model, and a gradient boosting model—using only the miRNAs selected by that method. Each model is then used to predict responses from the test set, and the smallest of the three prediction errors is recorded. We implement all three models so that no method has an inherent advantage over another. For example, the miRNAs selected by IPSSL may be better suited to minimizing error in a linear model than those selected by IPSSGB, while the miRNAs selected by IPSSGB may be better suited to minimizing error in a gradient boosting model than those selected by IPSSL. The linear and random forest predictive models are implemented with scikit-learn (Pedregosa et al., 2011) and gradient boosting with XGBoost (Chen and Guestrin, 2016), always with default parameters. For continuous responses, in each CV step we subtract the mean of the training responses from all responses, training and test, and scale all responses by the empirical standard deviation of the training responses.
Table 2 shows the top five miRNAs related to prognosis for each IPSS method, as ranked by their median $q$-values over the 10 CV steps. In Section S3, we further investigate the findings in Table 2 via a literature review. Notably, we find numerous studies that link several of the top miRNAs selected by IPSSGB and IPSSRF, namely miR-30d, miR-93, miR-96 and miR-150, to ovarian cancer prognosis. Among these, IPSSL only ranks miR-96 in its top five, and we found little literature identifying its top-ranked feature, miR-1-2, to ovarian cancer. The results of the second study, identifying miRNAs that are related to miR-150, are shown in Figure 6. Here, IPSSGB achieves the lowest average prediction error over the 10 cross-validation steps of any method, while using just 15 to 20 miRNAs, achieving lower prediction error than even the method that uses all 587 miRNAs to predict miR-150 levels (the dashed gray line in Figure 6). Figure 7 is similar to Figure 6, except now the response is tumor purity, the proportion of cancerous cells in a tissue sample. Here, we see IPSSGB performs best with the fewest features, while the top 30 features selected by IPSSRF achieve essentially identical predictive performance to the method that uses all 588 miRNAs.
IPSSGB IPSSRF IPSSL
miR-96(0.2) miR-1270(0.27) miR-1-2(0.07)
miR-150(0.28) miR-93(0.36) miR-1270(0.08)
miR-1270(0.43) miR-30d(0.36) miR-96(0.09)
miR-30d(0.43) miR-342(0.36) miR-342(0.09)
miR-1301(0.46) miR-150(0.41) miR-934(0.12)
Table 2. MicroRNA and prognosis. The top five miRNAs related to prognosis according to each IPSS method, as ranked by their median $q$-values over the 10 cross-validation steps.
Figure 6. MicroRNA and miR-150. The horizontal and vertical axes show the average number of features selected and the average prediction error over the 10 cross-validation steps, respectively. Curves for the IPSS methods and SSBoost are obtained by varying the target E(FP) between 1/8 and 10. Points for these methods correspond to target E(FP)s of 1/8, 1/4, 1/2, 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, with larger target E(FP)s generally corresponding to more features selected and lower prediction error. Similarly, points for KOBT correspond to target FDRs of 0.01, 0.025, 0.05, 0.1, 0.15, 0.2, 0.25, 0.3, and 0.4. The points for RF and XGBoost correspond to selecting between 1 and 30 features (hence, the number of features selected by these two methods is deterministic). Boruta and Vita do not have error control parameters and are therefore indicated by single points rather than curves.
Figure 7. MicroRNA and tumor purity. This is the same as Figure 6, except now the response is tumor purity, the proportion of cancerous cells in a sample. We again see the IPSS methods outperform other methods, achieving similar or better predictive performance with relatively few features.
6.  Discussion
We have demonstrated that IPSSRF and IPSSGB achieve superior results in terms of false positive control, detecting true positives, and computation time. More broadly, IPSS for thresholding is a general framework whose theory and implementation apply to arbitrary importance scores. For instance, examples of other scores include $p$-values (with smaller $p$-values indicating greater importance), Shapley values (used to quantify the contribution of individual features to neural networks and other machine learning models), and loadings in principal components analysis (which quantify the contribution of each feature to a given principal component). The main practical limitation to consider is the cost of computing the relevant importance scores, since IPSS must compute these scores on many random subsamples of the data.
We have also introduced efp scores and shown that, in addition to controlling E(FP), they can control the FDR and estimate $q$-values. Storey (2003) showed that $q$-values admit a Bayesian interpretation, suggesting a link between IPSS and Bayesian feature selection that could be an interesting line of future work. It would also be interesting to study whether $q$-values estimated from efp scores are more accurate than $q$-values estimated from other quantities, such as $p$-values (Storey and Tibshirani, 2003).
Finally, a more ambitious goal is to extend IPSS to unsupervised feature selection problems (that is, feature selection when there is no response variable) and non-iid data. The PCA-based scores mentioned above provide at least one way to apply IPSS in an unsupervised setting. Developing a rigorous approach to IPSS for non-iid data could provide novel methods for nonparametric feature selection with error control for time series, networks, and longitudinal data.
Data availability and code
All of the code and data used in this work are freely available at https://github.com/omelikechi/ipss. A Python package for implementing integrated path stability selection, including IPSSGB and IPSSRF, is freely available at https://pypi.org/project/ipss/
Acknowledgements
D.B.D. and O.M. were supported in part by funding from Merck & Co. and the National Institutes of Health (NIH) grant R01ES035625. J.W.M. and O.M. were supported in part by the Collaborative Center for X-linked Dystonia Parkinsonism (CCXDP). J.W.M. was supported in part by the National Institutes of Health (NIH) grant R01CA240299.
Supplementary material for “Fast nonparametric feature selection with error control using integrated path stability selection”
S1.  Additional IPSS details
We provide an algorithm that implements integrated path stability selection (IPSS) for thresholding (Section S1.1), elaborate on the theory of IPSS and its connection to efp scores (Section S1.2), and describe and study the IPSS parameters in greater detail (Section S1.3).
S1.1.  Algorithm.
Algorithm S1, discussed in Section 2.3, implements IPSS for thresholding. The number of grid points used to evaluate the integrals in Algorithm S1 is always $K=100$. Like many of the other IPSS parameters, $K$ is inconsequential provided it is sufficiently large; in our experience, values greater than 25 suffice. This is because the function $f(x)=(2x-1)^{3}\operatorname{\mathds{1}}(x\geq 0.5)$, the paths $\lambda\mapsto\hat{\pi}_{j}(\lambda)$ (which are monotonically increasing functions of $\lambda$), the quantity $\mathcal{I}(\Lambda)$, and the measure $\mu$ are all very numerically stable.
Algorithm S1 Integrated path stability selection for thresholding
1:Data $Z_{1:n}$, importance function $\Phi$, number of grid points $K$, and number of iterations $B$.
2:(Optional) Preselect features, as described in Section S1.4.
3:for $b=1,\ldots,B$ do
4:     Randomly select $A_{2b-1},A_{2b}\subseteq\{1,\ldots,n\}$ with $A_{2b-1}\cap A_{2b}=\varnothing$ and $\lvert A_{2b-1}\rvert=\lvert A_{2b}\rvert=\lfloor n/2\rfloor$.
5:     Evaluate $\Phi_{Z_{A_{2b-1}}}(j)$ and $\Phi_{Z_{A_{2b}}}(j)$ for $j=1,\dots,p$.
6:end for
[…]
8:Define a $\lambda$ grid with upper bound $\lambda_{\mathrm{max}}$, e.g., $\lambda_{\mathrm{max}}=\lambda_{0}>\lambda_{1}>\cdots>\lambda_{K}=\lambda_{\mathrm{max}}/10^{8}$.
9:Initialize $\lambda_{\mathrm{min}}\leftarrow\lambda_{\mathrm{max}}$ and $k\leftarrow 0$.
10:while $\mathcal{I}([\lambda_{\mathrm{min}},\lambda_{\mathrm{max}}])<0.05$ do
11:     $\hat{S}_{\lambda}(Z_{A_{b}})=\{j:\Phi_{Z_{A_{b}}}(j)\geq\lambda\}$ for $b=1,\dots,2B$.
12:     $\lambda_{\mathrm{min}}\leftarrow\lambda_{k+1}$ followed by $k\leftarrow k+1$.
13:end while
14:$\Lambda\leftarrow[\lambda_{\mathrm{min}},\lambda_{\mathrm{max}}]$.
15:Evaluate estimated selection probability $\hat{\pi}_{j}(\lambda)=\frac{1}{2B}\sum_{b=1}^{2B}\operatorname{\mathds{1}}\!\big{(}j\in\hat{S}_{\lambda}(Z_{A_{b}})\big{)}$ for $j=1,\dots,p$.
16:Evaluate the integral ${\int}_{\Lambda}f(\hat{\pi}_{j}(\lambda))\mu(d\lambda)$ for $j=1,\dots,p$.
17:$\mathtt{efp}_{Z_{1:n}}(j)=\mathcal{I}(\Lambda)/{\int}_{\Lambda}f(\hat{\pi}_{j}(\lambda))\mu(d\lambda)$ for $j=1,\dots,p$.
S1.2.  Theory: From IPSS to efp scores.
Given the estimated selection probabilities $\hat{\pi}_{j}$, an interval $\Lambda\subseteq(0,\infty)$, a probability measure $\mu$ on $\Lambda$, a function $f:[0,1]\to\mathbb{R}$, and a threshold $\tau$, Melikechi and Miller (2024) define the set of features selected by IPSS by
$\hat{S}_{\mathrm{IPSS}}(\tau)$ =$=\Big{\{}j:{\int}_{\Lambda}f(\hat{\pi}_{j}(\lambda))\mu(d\lambda)\geq\tau\Big{\}}.$ (S1.1)
The integral incorporates information about the selection probabilities over all of $\Lambda$, eliminating the need to select features based on individual $\lambda$ values. The function $f$ transforms the selection probabilities for better performance. Only certain choices of $f$ are known to yield valid efp scores. Melikechi and Miller (2024) prove the following result (see Theorem 4.1 therein). Let $q(\lambda)=\mathrm{E}\lvert\hat{S}_{\lambda}(Z_{1:\lfloor n/2\rfloor})\rvert$ be the expected number of features selected by $\hat{S}_{\lambda}$ on half the data and, for $m\in\mathbb{N}$, define $h_{m}(x)=(2x-1)^{m}\operatorname{\mathds{1}}(x\geq 0.5)$. For any $\Lambda$ and $\mu$, if
$\max_{j\in S^{c}}\,\mathbb{P}\bigg{(}j\in\bigcap_{b=1}^{m^{\prime}}\big{(}\hat{S}_{\lambda}(Z_{A_{2b-1}})\cap\hat{S}_{\lambda}(Z_{A_{2b}})\big{)}\bigg{)}$ $\leq(q(\lambda)/p)^{2m^{\prime}},$ (S1.2)
for all $\lambda\in\Lambda$ and $m^{\prime}\in\{1,\dots,m\}$, then IPSS with $f=h_{m}$ satisfies
$\mathrm{E}\lvert\hat{S}_{\mathrm{IPSS}}(\tau)\cap S^{c}\rvert$ $\leq\frac{\mathcal{I}_{m}(\Lambda)}{\tau}$ (S1.3)
for a constant $\mathcal{I}_{m}(\Lambda)$ whose explicit form is given in Theorem 4.1 in Melikechi and Miller (2024). Thus, setting $\mathtt{efp}_{Z_{1:n}}(j)=\mathcal{I}_{m}(\Lambda)/\int_{\Lambda}h_{m}(\hat{\pi}_{j}(\lambda))\mu(d\lambda)$ and $\hat{S}(t)=\{j:\mathtt{efp}_{Z_{1:n}}(j)\leq t\}$, we have
$\hat{S}(t)$ =$=\bigg{\{}j:\frac{\mathcal{I}_{m}(\Lambda)}{\int_{\Lambda}h_{m}(\hat{\pi}_{j}(\lambda))\mu(d\lambda)}\leq t\bigg{\}}=\bigg{\{}j:{\int}_{\Lambda}h_{m}(\hat{\pi}_{j}(\lambda))\mu(d\lambda)\geq\frac{\mathcal{I}_{m}(\Lambda)}{t}\bigg{\}}=\hat{S}_{\mathrm{IPSS}}\bigg{(}\frac{\mathcal{I}_{m}(\Lambda)}{t}\bigg{)}$
and hence
$\mathrm{E}(\mathrm{FP}(t))$ $=\mathrm{E}\lvert\hat{S}(t)\cap S^{c}\rvert=\mathrm{E}\big{\lvert}\hat{S}_{\mathrm{IPSS}}\big{(}\tfrac{\mathcal{I}_{m}(\Lambda)}{t}\big{)}\cap S^{c}\big{\rvert}\leq t,$
where the inequality holds by Equation S1.3.
The above derivation shows how valid efp scores for IPSS with $f=h_{m}$ are obtained under the conditions of Theorem 4.1 in Melikechi and Miller (2024). In particular, we see that by defining efp scores as above, the set $\hat{S}(t)=\{j:\mathtt{efp}_{Z_{1:n}}(j)\leq t\}$ is identical to $\hat{S}_{\mathrm{IPSS}}(\tau)$ when $\tau=\mathcal{I}_{m}(\Lambda)/t$. The main condition of the theorem, Equation S1.2, upper bounds the maximum probability that an unimportant feature is simultaneously selected on both halves of the data, $Z_{A_{2b-1}}$ and $Z_{A_{2b}}$, in $m^{\prime}$ independent tries; see Melikechi and Miller (2024) for details. The following result, part of Theorem 4.2 in Melikechi and Miller (2024), gives the form of $\mathcal{I}_{3}(\Lambda)$, which corresponds to the function $f=h_{3}$ that is used for IPSS throughout the main text.
Theorem S1.1.
Let $\mu$ be a probability measure on $\Lambda\subseteq(0,\infty)$, let $\tau\in(0,1]$, and define $\hat{S}_{\mathrm{IPSS}}(\tau)$ as in Equation S1.1 with $f=h_{3}$. If Equation S1.2 holds for all $\lambda\in\Lambda$ and $m^{\prime}\in\{1,2,3\}$, then
$\mathrm{E}(\mathrm{FP}(\tau))$ $\leq\frac{1}{\tau}\int_{\Lambda}\bigg{(}\frac{q(\lambda)^{2}}{B^{2}p}+\frac{3q(\lambda)^{4}}{Bp^{3}}+\frac{q(\lambda)^{6}}{p^{5}}\bigg{)}\mu(d\lambda),$ (S1.4)
where $\mathrm{E}(\mathrm{FP}(\tau))=\mathrm{E}\lvert\hat{S}_{\mathrm{IPSS}}(\tau)\cap S^{c}\rvert$ is the expected number of false positives selected by IPSS.
For some intuition about the bound in Equation S1.4, observe that taking $B\to\infty$ yields $\mathrm{E}(\mathrm{FP}(\tau))\leq\tau^{-1}\int_{\Lambda}\big{(}q(\lambda)^{6}/p^{5}\big{)}\mu(d\lambda)$. In comparison, other versions of stability selection upper bound $\mathrm{E}(\mathrm{FP}(\tau))$ by $q(\lambda)^{2}/p$ (Meinshausen and Bühlmann, 2010; Shah and Samworth, 2013), which is orders of magnitude larger than $q(\lambda)^{6}/p^{5}$ when $q(\lambda)\ll p$, as is often the case for most values of $\lambda$ in $\Lambda$. This and the contribution of $B$ in the denominators in Equation S1.4 largely explain the strength of the IPSS bound in Theorem S1.1 relative to previous bounds, and hence the tightness of the efp scores of IPSS with $f=h_{3}$ relative to the efp scores of other versions of stability selection. Further theoretical and empirical comparisons between Equation S1.4 and other stability selection bounds are available in Melikechi and Miller (2024).
S1.3.  Parameters.
As noted above, the function $f(x)=(2x-1)^{3}\operatorname{\mathds{1}}(x\geq 0.5)$ used when implementing IPSS in the main text is determined by the availability and strength of the theoretical bound in Theorem S1.1. Similarly, the choice of $\lfloor n/2\rfloor$ samples used to construct the selection probabilities $\hat{\pi}_{j}(\lambda)$ are required for stability selection theorems (not just Theorem S1.1) to hold, and it is unclear how to adapt their proofs to accommodate other sample sizes. Thus, $f$ and $\lfloor n/2\rfloor$ are theoretical necessities rather than free parameters.
The interval $\Lambda=[\lambda_{\mathrm{min}},\lambda_{\mathrm{max}}]$ is determined by setting $\lambda_{\mathrm{max}}$ large enough that no features are selected (see, for example, 7 in Algorithm S1), and setting $\lambda_{\mathrm{min}}$ such that the integral
$\mathcal{I}(\Lambda)$ $=\int_{\Lambda}\bigg{(}\frac{q(\lambda)^{2}}{B^{2}p}+\frac{3q(\lambda)^{4}}{Bp^{3}}+\frac{q(\lambda)^{6}}{p^{5}}\bigg{)}\mu(d\lambda),$
in Equation S1.4 is equal to a fixed cutoff $C$. As noted in the main text, we always use $C=0.05$, but results are largely independent of this choice (Figures S1 and S2). Figures S3 and S4 show IPSS is also indifferent to $\mu$. Here, the parameter $\delta$ determines the measure $\mu_{\delta}(d\lambda)=z_{\delta}^{-1}\lambda^{-\delta}d\lambda$, where the normalizing constant $z_{\delta}$ is easily computed in closed form (Melikechi and Miller, 2024). The probability measure that we always use in the main text, $\mu_{1}$, averages over $\Lambda$ on a log scale, while $\mu_{0}$ averages on a linear scale. The insignificance of $C$ and $\mu$ is unsurprising: Intuitively, the efp scores for IPSS depend primarily on $f$ and the integrand in $\mathcal{I}(\Lambda)$. The actual value of $\mathcal{I}(\Lambda)$ is much less important since the efp scores depend on the relative quantities $\mathcal{I}(\Lambda)/\int_{\Lambda}f(\hat{\pi}_{j}(\lambda))\mu(d\lambda)$ rather than on $\mathcal{I}(\Lambda)$ itself. Since $C$ and $\mu$ only affect the value of the bound, not the integrand, they contribute little to the actual performance of IPSS.
Figure S1. Dependence on cutoff (regression). IPSSGB is applied to regression data simulated as in Section 4 for cutoff values $C\in\{0.025,0.05,0.075,0.1,0.15\}$, averaged over 100 trials. There is little difference between the average true and false positive values for different choices of $C$ over a wide range of target E(FP) values. Figures S2, S3, and S4 show similar results for classification and dependence on $\mu$.
Figure S2. Dependence on cutoff (classification). IPSSGB is applied to classification data simulated as in Section 4 for cutoff values $C\in\{0.025,0.05,0.075,0.1,0.15\}$, averaged over 100 trials. There is little difference between the average true and false positive values for different choices of $C$ over a wide range of target E(FP) values.
Figure S3. Dependence on $\mu$ (regression). IPSSGB is applied to regression data simulated as in Section 4 for cutoff values $\delta\in\{0,0.5,1,1.5,2\}$, averaged over 100 trials, where $\delta$ determines to the measure $\mu_{\delta}(d\lambda)=z_{\delta}^{-1}\lambda^{-\delta}d\lambda$. There is little difference between the average true and false positive values for different choices of $\delta$ over a wide range of target E(FP) values.
Figure S4. Dependence on $\mu$ (classification). IPSSGB is applied to classification data simulated as in Section 4 for cutoff values $\delta\in\{0,0.5,1,1.5,2\}$, averaged over 100 trials, where $\delta$ determines to the measure $\mu_{\delta}(d\lambda)=z_{\delta}^{-1}\lambda^{-\delta}d\lambda$. There is little difference between the average true and false positive values for different choices of $\delta$ over a wide range of target E(FP) values.
S1.4.  Preselection.
Many feature selection algorithms, including most in Section 3, employ some form of screening, or preselection, as an initial step in the selection process. This can be especially helpful in high dimensions for reducing noise and runtime. When $p>200$, we preselect features for IPSSGB, IPSSRF, and the other stability selection-based methods (IPSSL and SSBoost, detailed below) by running their respective selection algorithms on the full dataset three times—that is, computing $\Phi_{Z_{1:n}}$ three times—and keeping only the $p_{\text{pre}}=\max\{200,p/20\}$ features with the largest average scores across all three trials. Note that $\Phi_{Z_{1:n}}$ is stochastic, given $Z_{1:n}$, due to randomness in the gradient boosting and random forest algorithms. This preselection step does not affect the theoretical control on E(FP) since $\lvert\hat{S}_{\mathrm{IPSS},\mathrm{pre}}\cap S^{c}\rvert=\lvert\hat{S}_{\mathrm{IPSS},\mathrm{pre}}\cap S^{c}_{\mathrm{pre}}\rvert$, where $\hat{S}_{\mathrm{IPSS},\mathrm{pre}}$ are the features selected by IPSS using only the preselected features, and $S^{c}_{\mathrm{pre}}$ are the preselected features in $S^{c}$. Of course, preselection risks throwing away important features, potentially increasing the number of false negatives. However, we have yet to observe a case in which an important feature that was not in the top $p_{\mathrm{pre}}$ preselected features was selected by IPSS when the full set of features was used instead. In other words, a feature that does not make the top $p_{\mathrm{pre}}$ is highly unlikely to be selected by IPSS regardless of whether or not preselection is used.
S2.  Simulation details and additional results
Algorithm S2 describes the procedure we use to generate data in the simulation study in Section 4. See also Figure 1 in the main text for a diagrammatic representation of this process. In all steps, “randomly select” means select a parameter uniformly at random from its domain, which are shown in Table S1. The randomized function $f_{\theta}:\mathbb{R}\to[-1,1]$ that links the features to the response is
$f_{\theta}(x)$ =$=\begin{cases}\frac{\delta_{1}}{2}\left(1+\tanh(\alpha(\delta_{2}x-\beta))\right)&\text{with probability}\ 1/2,\\ \delta_{1}\exp(-\gamma x^{2})&\text{with probability}\ 1/2,\end{cases}$ (S2.1)
where each component of $\theta=(\alpha,\beta,\gamma,\delta_{1},\delta_{2})$ is drawn uniformly at random prior to each trial according to Table S1. The values of $\alpha$ and $\gamma$ determine steepness of the curves, $\beta$ shifts $\tanh$ horizontally, and $\delta_{1}$ and $\delta_{2}$ reflect the functions about the horizontal and vertical axes, respectively. Figure S7 shows five realizations of $f_{\theta}$, illustrating the many ways it can influence the response, $y$. For example, if the function in the left-most panel of Figure S7 is applied to the genes in a given partition of $S$, then those genes will only significantly affect $y$ if their collective expression level is positive, while collective expression levels less than $-1$ will have virtually no effect on $y$.
Parameter $\alpha$ $\beta$ $\gamma$ $\delta_{1}$ $\delta_{2}$ $p_{\mathrm{true}}$ SNR $u$
Range $(0.5,1.5)$ $(-1,1)$ $(1,3)$$\{-1,1\}$$\{-1,1\}$$\{10,\dots,30\}$ $(0.5,2)$ $(1,3)$
Table S1. Simulation parameters. Parameters are drawn uniformly at random from their corresponding ranges prior to each simulation trial.
Algorithm S2 Data generation for simulation study (one trial)
1:RNA-seq data $X_{\mathrm{full}}\in\mathbb{R}^{596\times 6426}$, number of samples $n$, number of features $p$, number of true features $p_{\mathrm{true}}$, signal-to-noise ratio SNR, function parameter domain $\Theta$.
2:Randomly select $n$ rows and $p$ columns of $X_{\mathrm{full}}$. Denote the resulting matrix by $X\in\mathbb{R}^{n\times p}$.
3:Standardize the columns of $X$ to have mean 0 and variance 1.
4:Randomly select true features $S\subseteq\{1,\dots,p\}$ with $\lvert S\rvert=p_{\mathrm{true}}$.
5:Randomly select $G\in\{\lfloor p_{\mathrm{true}}/2\rfloor,\dots,p_{\mathrm{true}}\}$. Partition $S$ into $G$ disjoint groups, $S=\bigsqcup_{g=1}^{G}S_{g}$.
6:Initialize the signal, $\eta\leftarrow(0,\ldots,0)^{\mathtt{T}}\in\mathbb{R}^{n}$.
7:for $g=1,\dots,G$ do
8:     $\xi_{g}\leftarrow\sum_{j\in S_{g}}X_{j}$ where $X_{j}\in\mathbb{R}^{n}$ is the $j$th column of $X$.
9:     Standardize $\xi_{g}$ to have mean 0 and variance 1.
10:     Randomly select a function parameter $\theta\in\Theta$.
11:     $\eta\leftarrow\eta+f_{\theta}(\xi_{g})$ with $f_{\theta}$ applied to $\xi_{g}\in\mathbb{R}^{n}$ componentwise.
12:end for
13:For regression: Draw $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$ with $\sigma^{2}=\sum_{i=1}^{n}\eta_{i}^{2}/(n\,\mathrm{SNR})$ and set $y_{i}\leftarrow\eta_{i}+\epsilon_{i}$.
14:For classification: Draw $u\sim\mathrm{Uniform}(1,3)$, then $y_{i}\sim\mathrm{Bernoulli}(\pi_{i})$ where $\pi_{i}=1/(1+\exp(-u\eta_{i}))$.
15:Features $X\in\mathbb{R}^{n\times p}$, responses $y\in\mathbb{R}^{n}$, and important features $S\subseteq\{1,\ldots,p\}$.
Figure S5. Regression simulation. Empirical E(FP) and E(TP) versus target E(FP) for each IPSS method. The dotted black line shows where E(FP) = target E(FP).
Figure S6. Classification simulation. Empirical E(FP) and E(TP) versus target E(FP) for each IPSS method. The dotted black line shows E(FP) = target E(FP).
Figure S7. Some realizations of the randomized function $f_{\theta}$.
One benefit of IPSS is that changing the target E(FP) does not require rerunning the algorithm. Figures S5 and S6 show E(TP) and E(FP) as functions of the target E(FP). IPSSGB has the best performance in terms of true positives across essentially all target E(FP) values. IPSSGB and IPSSRF perform well in the regression setting, and both slightly overshoot the target E(FP) in classification for values less than 6. IPSSL performs poorly in general, having relatively few true positives on average and false positives that exceed the target E(FP).
S3.  Additional ovarian cancer results
In Section S3.1, we expand upon the ovarian cancer results presented in Section 5 of the main text, namely identifying miRNAs related to prognosis, and miRNAs related to miR-150, respectively. In Section S3.2, we perform a similar analysis but for RNA-seq data from the same cohort, identifying genes related to ovarian cancer prognosis, and genes related to one of the genes, AKT2, that was found to be related to prognosis.
S3.1.  MicroRNA and ovarian cancer.
We briefly summarize results from the literature that link the miRNAs in Table 2 to ovarian cancer, with a specific focus on prognosis. For many of these miRNAs, there are numerous studies connecting them to ovarian cancer beyond what we present. Thus, to roughly quantify the relevance of the miRNAs in Table 2 to ovarian cancer, we searched the specific miRNA together with “ovarian cancer” in the Web of Science database, and report both the number of publications and the total number of citations across all returned results. We emphasize that these metrics are less important than the papers themselves. For example, citation counts favor older publications, and our search criterion does not guarantee that the returned results are relevant to the problem at hand (though the summaries below show that at least some of them are).
Table S2 is similar to Table 2, except that it shows the total number of publications and citations from our Web of Science literature search and includes some $q$-values that were omitted from Table 2, since Table 2 only reported the five miRNA with the lowest $q$-values for each method. Dashed lines indicate that the feature was not identified by the corresponding method at all.
miRNA Citations Publications IPSSGB IPSSRF IPSSL
miR-150 1635 26 0.28 0.41 0.2
miR-93 1270 28 0.36
miR-30d 596 12 0.43 0.36 0.16
miR-96 338 10 0.2 0.53 0.09
miR-342 201 9 0.36 0.09
miR-1270 175 5 0.43 0.27 0.08
miR-934 27 2 0.12
miR-1301 11 3 0.46
miR-1-2 5 1 0.07
Table S2. MicroRNA and prognosis. MiRNA are ordered by citations. A missing $q$-value means that miRNA was not selected by the corresponding version of IPSS.
miR-1-2. Kandettu et al. (2022) found that miR-1-2 is differentially expressed between cancerous and non-cancerous ovarian cancer cells, but we found no literature linking miR-1-2 to prognosis.
miR-30d. Ye et al. (2015) found that miR-30d suppresses ovarian cancer progression by reducing the levels of Snail, a protein involved in making cancer cells more invasive. They concluded that miR-30d could be used as a treatment for ovarian cancer. Lee et al. (2012) found that miR-30d is associated with “significantly better disease-free or overall survival” in ovarian cancer patients.
miR-93. Fu et al. (2012) found that miR-93 is significantly upregulated in ovarian cancer cells that are resistant to the chemotherapy drug cisplatin. They also found that miR-93 targets the tumor suppressor gene PTEN and plays a role in the AKT signaling pathway. They concluded that further study of miR-96 may yield therapeutic strategies for overcoming cisplatin-resistant ovarian cancer cells. Meng et al. (2015) found that miR-93 is a potential biomarker of ovarian cancer.
miR-96. Liu et al. (2019) found that overexpression of miR-96 promotes cell proliferation and migration in ovarian cancer cells. They conclude that targeting miR-96 is a potentially promising strategy for treating ovarian cancer. They also report that miR-96 inhibits phosphorylation of AKT, a gene identified by IPSSGB as being relevant to ovarian cancer prognosis (Section S3.2). Yang et al. (2020) found that individuals with low-levels of miR-96 “suffered more advanced tumor staging and a worse overall survival” and also identified miR-96 as a potential therapeutic target.
miR-150. Jin et al. (2014) found significant associations between miR-150 downregulation and “aggressive clinicopathological features” in ovarian cancer patients, as well as reduced overall and progression-free survival. They also identified miR-150 expression as a prognostic biomarker in ovarian cancer. Kim et al. (2017) found that downregulation of miR-150 is associated with resistance to paclitaxel, a chemotherapy drug used to treat ovarian cancer. They also report that treatment with pre-miR-150 resensitized cancer cells to paclitaxel, making the drug more effective.
miR-342. Dou et al. (2020) found that miR-342 inhibits the proliferation, invasion, and migration of ovarian cancer cells, and promotes the death of these cells. The study also showed that miR-342 decreases the expression of key proteins involved in the Wnt/$\beta$-catenin signaling pathway, which may explain its effects on reducing ovarian cancer cell viability and growth.
miR-934. Hu et al. (2019) found that miR-934 promotes ovarian cancer cell proliferation and inhibits the death of ovarian cancer cells by targeting the gene BRMS1L, making it is a potential biomarker or therapeutic target for ovarian cancer.
miR-1270. Ghafouri-Fard et al. (2022) found that miR-1270 plays a role in sensitivity to the chemotherapy drug cisplatin.
miR-1301. Yu and Gao (2020) found that targeting miR-1301 can inhibit the proliferation of cells that are resistant to the chemotherapy drug cisplatin, thus reducing the occurrence and development of drug-resistant ovarian cancer.
Tables S3 and S4 show the top ten features for each IPSS method, ranked by their median $q$-values over the 10 CV steps, when the features are miRNA and the responses are miR-150 (Table S3) and tumor purity (Table S4), respectively. Tumor purity is the proportion of cancerous cells in a sample. Observe that the $q$-values in Table S3 are much smaller than those in Table S2. This is to be expected since the signal between miRNAs, in particular, between miR-150 and certain other miRNAs, is likely much stronger than the signal between miRNAs and prognosis.
IPSSGB IPSSRF IPSSL
miR-142(0.031) miR-146a(0.016) miR-146a(0.028)
miR-146a(0.031) miR-142(0.016) miR-142(0.028)
miR-1247(0.031) miR-223(0.016) miR-1247(0.028)
miR-642(0.032) miR-642(0.016) miR-139(0.028)
miR-342(0.036) miR-155(0.016) miR-342(0.029)
miR-139(0.042) miR-1247(0.016) miR-642(0.029)
miR-1249(0.051) miR-766(0.016) miR-96(0.029)
miR-96(0.064) miR-342(0.017) miR-511-1(0.03)
miR-204(0.072) miR-1228(0.018) miR-34b(0.032)
miR-34b(0.074) miR-21(0.02) miR-184(0.033)
Table S3. MicroRNA and miR-150. Top ranked miRNAs for each method in terms of their $q$-values, shown in parentheses.
IPSSGB IPSSRF IPSSL
miR-150(0.05) miR-150(0.02) miR-150(0.05)
miR-22(0.05) miR-22(0.02) miR-22(0.05)
miR-140(0.08) miR-140(0.02) miR-142(0.05)
miR-142(0.08) miR-142(0.02) miR-140(0.05)
miR-1224(0.12) miR-146a(0.03) miR-25(0.06)
miR-103-2(0.16) miR-223(0.03) miR-1224(0.09)
miR-223(0.23) miR-214(0.04) miR-135a-2(0.1)
miR-181c(0.36) miR-132(0.04) miR-155(0.11)
miR-155(0.49) miR-642(0.05) miR-1247(0.12)
miR-25(0.52) miR-199b(0.05) miR-34c(0.13)
Table S4. MicroRNA and tumor purity. Top ranked miRNAs for each method in terms of their $q$-values, shown in parentheses.
S3.2.  RNA-seq and ovarian cancer.
In this section, we perform a similar analysis to the miRNA study in Sections 5 and S3.1, but with the 6426 genes from the RNA-seq data described in Section 4.
Gene Citations Publications IPSSGB IPSSRF IPSSL
AKT2 9088 118 0.25
CD38 852 30 0.11 0.31
PPL 151 8 0.8 0.26
HOXB4 127 7 0.36
ZFHX4 105 7 0.78 0.17
SLAMF7 101 2 0.38 0.36
WNK1 96 3 0.42 0.23
SLAMF1 58 5 0.38
PARD6B 61 1 0.28 0.33
UGT8 37 1 0.17 0.41
AP2A1 37 2 0.4
PROSC 0 0 0.33 0.46 0.2
MRAP2 0 0 0.09 0.39 0.78
OR51A4 0 0 0.59 0.2
CCDC84 0 0 0.2
SH3BGR 0 0 0.07
Table S5. RNA-seq and prognosis. Genes are ordered by citations. A missing $q$-value means that gene was not selected by the corresponding version of IPSS. *A literature search revealed that the gene MRAP2 (also called C6orf117), while not directly related to ovarian cancer, is very much related to obesity, which is linked to poor prognosis and reduced survival times in cancer patients (Chan et al., 2009).
Figure S8. RNA-seq and prognosis. No method does well at predicting prognosis relative to the full dataset. This is unsurprising, given the many factors that can contribute to poor outcomes which are unlikely to be captured by RNA-seq data alone. Nevertheless, several of the genes identified by the various feature selection methods do appear to be related to ovarian cancer prognosis; see Table S5.
Figure S9. RNA-seq and AKT2. As in several of the other studies, IPSSGB achieves superior predictive performance with fewer features compared to other methods.
IPSSGB IPSSRF IPSSL
SUPT5H(0.01) SUPT5H(0.01) SUPT5H(0.02)
EGLN2(0.01) PAK4(0.01) EGLN2(0.02)
PAK4(0.01) ADCK4(0.01) ADCK4(0.02)
ADCK4(0.01) EGLN2(0.01) BLVRB(0.02)
BLVRB(0.01) ZNF574(0.01) SPSB3(0.03)
ITPKC(0.02) GSK3A(0.01) HKR1(0.03)
HKR1(0.04) SARS2(0.01) CAV2(0.04)
ZMYND17(0.14) ITPKC(0.01) MAP3K10(0.07)
U2AF2(0.19) CLPTM1(0.01) ACRC(0.08)
SPSB3(0.2) ZNF526(0.01) ITPKC(0.09)
Table S6. RNA-seq and AKT2. Top ranked genes for each method in terms of their $q$-values, shown in parentheses.

### v2

@@
-Fast nonparametric feature selection with error control using integrated path stability selection
+Nonparametric IPSS Fast, flexible feature selection with false discovery control
 Abstract.
-Feature selection can greatly improve performance and interpretability in machine learning problems.
-However, existing nonparametric feature selection methods either lack theoretical error control or fail to accurately control errors in practice.
-Many methods are also slow, especially in high dimensions.
-In this paper, we introduce a general feature selection method that applies integrated path stability selection to thresholding to control false positives and the false discovery rate.
-The method also estimates $q$-values, which are better suited to high-dimensional data than $p$-values.
-We focus on two special cases of the general method based on gradient boosting (IPSSGB) and random forests (IPSSRF).
-Extensive simulations with RNA sequencing data show that IPSSGB and IPSSRF have better error control, detect more true positives, and are faster than existing methods.
-We also use both methods to detect microRNAs and genes related to ovarian cancer, finding that they make better predictions with fewer features than other methods.
+Feature selection is a critical task in machine learning and statistics.
+However, existing feature selection methods either (i) rely on parametric methods such as linear or generalized linear models, (ii) lack theoretical false discovery control, or (iii) identify few true positives.
+Here, we introduce a general feature selection method with finite-sample false discovery control based on applying integrated path stability selection (IPSS) to arbitrary feature importance scores.
+The method is nonparametric whenever the importance scores are nonparametric, and it estimates $q$-values, which are better suited to high-dimensional data than $p$-values.
+We focus on two special cases using importance scores from gradient boosting (IPSSGB) and random forests (IPSSRF).
+Extensive nonlinear simulations with RNA sequencing data show that both methods accurately control the false discovery rate and detect more true positives than existing methods.
+Both methods are also efficient, running in under 20 seconds when there are 500 samples and 5000 features.
+We apply IPSSGB and IPSSRF to detect microRNAs and genes related to cancer, finding that they yield better predictions with fewer features than existing approaches.
 1.  Introduction
-Identifying the important features in a dataset can greatly improve performance and interpretability in machine learning problems (Theng and Bhoyar, 2024).
+Identifying the important features in a dataset can greatly improve performance and interpretability in machine learning and statistical problems (Theng and Bhoyar, 2024).
 For example, in genomics, often only a small fraction of genes (features) are related to a disease of interest (response).
 By identifying these genes, scientists can save time and resources while gaining insights that would be difficult to uncover otherwise (Theng and Bhoyar, 2024).
-A common form of feature selection is thresholding: Given a function $\Phi$ that assigns a degree of importance $\Phi(j)$ to each feature $j$ in a dataset, one seeks a value $\lambda$ such that the selected set $\hat{S}_{\lambda}=\{j:\Phi(j)\geq\lambda\}$ contains important features and its complement, $\hat{S}^{c}_{\lambda}=\{j:\Phi(j)<\lambda\}$, does not.
-Choosing $\lambda$ can be challenging, and many methods for doing so come with few theoretical guarantees (Huynh-Thu et al., 2012).
-Stability selection is a popular technique for improving the performance of feature selection algorithms (Meinshausen and Bühlmann, 2010).
-However, its implementation relies on relatively weak upper bounds on the expected number of false positives, E(FP), that is, on the expected number of unimportant features that are selected.
-Recently, Melikechi and Miller (2024) proved that much stronger bounds hold for integrated path stability selection (IPSS), which consequently identifies more true positives in practice than stability selection.
-Until now, IPSS has only been applied to parametric linear models, and its generalization to more complex models has not been explored.
-In this work, we extend IPSS to nonlinear, nonparametric models by applying it to thresholding.
-The result is a feature selection method with tight upper bounds on E(FP) that are characterized by novel quantities called efp scores.
-In addition to controlling E(FP), efp scores also approximately control the false discovery rate (FDR) and estimate the $q$-values for each feature, which are more reliable and robust than $p$-values in genomics and other high-dimensional settings (Storey and Tibshirani, 2003).
-Lastly, by integrating over threshold values, IPSS for thresholding replaces the problem of specifying a single threshold $\lambda$ with the simpler task of specifying either a target E(FP) or a target FDR.
-We develop two specific instances of the general method: IPSS for gradient boosting (IPSSGB) and IPSS for random forests (IPSSRF).
-In simulations, we find that IPSSGB has better false positive control and identifies more true positives than nine other feature selection methods, and that both IPSSGB and IPSSRF outperform IPSS for linear models when the linearity assumptions of the latter are violated.
-More generally, IPSSGB and IPSSRF control E(FP) and FDR more accurately than other stability selection and knockoff-based methods with theoretical error control.
-In Section 5, we find that IPSSGB and IPSSRF effectively identify microRNAs and genes related to ovarian cancer, achieving better predictive performance than other feature selection methods while using fewer features.
-Both also run faster than leading methods.
+The goal of feature selection is to maximize the number of important features selected (true positives)—informally referred to here as power—while minimizing the number of unimportant features selected (false positives).
+In simulations, we find that popular methods without theoretical false discovery control, such as recursive feature elimination and Boruta (Kursa et al., 2010), often select many false positives (Section 3).
+Meanwhile, popular methods with false discovery control, namely stability selection (Meinshausen and Bühlmann, 2010; Shah and Samworth, 2013) and model-X knockoffs (Candes et al., 2018), have low power in simulations and select few features in practice (Sections 3 and 4).
+One reason for stability selection’s low power is its relatively weak theoretical upper bounds on the expected number of false positives, E(FP).
+Recently, Melikechi and Miller (2024) proved that much stronger bounds hold for integrated path stability selection (IPSS), which consequently identifies more true positives.
+Until now, IPSS has only been applied to generalized linear models, limiting its applicability since parametric assumptions are often violated or difficult to verify in practice.
+Thus, there is a need for nonparametric feature selection methods with theoretical false discovery control and greater power.
+In this work, we address this need by applying IPSS to arbitrary feature importance scores.
+The result is a general feature selection method with tight upper bounds on E(FP) that are characterized by novel quantities called efp scores.
+In addition to controlling E(FP), efp scores approximately control the false discovery rate (FDR) and estimate $q$-values for each feature, which are more reliable than $p$-values in genomics and other high-dimensional settings (Storey and Tibshirani, 2003).
+Our proposed method is nonparametric whenever the feature importance scores come from nonparametric models.
+We develop two specific instances of this: IPSS for gradient boosting (IPSSGB) and IPSS for random forests (IPSSRF).
+Like knockoffs, neither IPSSGB nor IPSSRF assume a specific functional relationship between the response and the features.
+Unlike knockoffs, neither method requires knowledge of the joint distribution of the features.
+In simulations, we find that IPSSGB and IPSSRF provide a better balance of FDR control and power than 12 other methods, and that IPSSGB performs best overall.
+In particular, both methods significantly outperform parametric versions of IPSS when the parametric assumptions are violated.
+In Section 4, we find that IPSSGB and IPSSRF successfully identify microRNAs and genes related to ovarian cancer and glioma, achieving better predictive performance than other feature selection methods while using fewer features.
+Both are also computationally efficient.
+Finally, another important aspect of feature selection is stability—the consistent selection of features across similar settings.
+Stability improves reproducibility, which is critical in many applications (Li et al., 2022).
+As their names suggest, stability selection and IPSS are designed to produce stable results by repeatedly applying a baseline feature selection algorithm to random subsamples of the data (Section 2.2).
+Nogueira et al.
+(2018) show that stability selection can produce significantly more stable results than its baseline algorithm.
+In this work, we focus on false discovery control and power; further study of the stability of IPSS, stability selection, and other stability-inspired methods like StabML-RFE (Li et al., 2022) is left to future work.
 Organization.
-In Section 2, we introduce efp scores, IPSS for thresholding in general, and IPSS for gradient boosting and random forests.
-In Section 3, we describe several leading feature selection methods, which are compared to IPSSGB and IPSSRF in a simulation study in Section 4.
-In Section 5, we analyze ovarian cancer data.
+In Section 2, we introduce efp scores, review IPSS, and present its extension to feature importance scores.
+In Section 3, we present our simulation studies, and in Section 4, we analyze ovarian cancer and glioma data.
 We conclude in Section 5 with a discussion.
 2.  Methods
-We introduce efp scores (Section 2.1), review IPSS (Section 2.2), introduce IPSS for thresholding (Section 2.3), and apply IPSS to gradient boosting and random forests (Section 2.4).
+In Section 2.1, we introduce efp scores.
+These quantities, assigned to each feature in the dataset, are used to perform feature selection with E(FP) control.
+In Section 2.2, we introduce IPSS, which is a general approach for constructing efp scores.
+In Section 2.3, we describe how IPSS can be applied to any feature importance score, focusing in particular on importance scores from tree-based methods such as gradient boosting and random forests.
 […]
-2.1.  efp scores.
-Suppose $S\subseteq\{1,\ldots,p\}$ is an unknown subset of true important features that we wish to estimate using $Z_{1:n}$.
+2.1  efp scores.
+Suppose $S\subseteq\{1,\ldots,p\}$ is an unknown subset of important features that we wish to estimate using $Z_{1:n}$.
 An efp (expected false positive) score is a function $\mathtt{efp}_{Z_{1:n}}:\{1,\dots,p\}\to[0,\infty)$ that depends on $Z_{1:n}$ and satisfies the following:
@@
 Informally, the false discovery rate is the expected ratio between the number of false positives and the total number of features selected, $\mathrm{FDR}=\mathrm{E(FP/(TP+FP))}$, and the $q$-value of feature $j$ is the smallest FDR when $j$ is selected (Storey, 2003).
-When $p$ is large, as is often the case in genomics, Storey (2003) showed that
+When $p$ is large, as is often the case in genomics, we have
 $\begin{split}\mathrm{pFDR}(t)&\approx\mathrm{FDR}(t)\approx\frac{\mathrm{E}(\mathrm{FP}(t))}{\mathrm{E}\lvert\hat{S}(t)\rvert}\leq\frac{t}{\mathrm{E}\lvert\hat{S}(t)\rvert},\end{split}$ (2.1)
-where $\mathrm{pFDR}(t)=\mathrm{E}(\mathrm{FP}(t)/\lvert\hat{S}(t)\rvert\mid\lvert\hat{S}(t)\rvert>0)$ is the positive false discovery rate, and the inequality holds by the definition of an efp score.
+where $\mathrm{pFDR}(t)=\mathrm{E}(\mathrm{FP}(t)/\lvert\hat{S}(t)\rvert\mid\lvert\hat{S}(t)\rvert>0)$ is the positive false discovery rate, the two approximations are from Storey (2003), and the inequality holds by the definition of an efp score.
 It follows that the $q$-value of feature $j$ satisfies
@@
 The largest such set is chosen to maximize true positives.
-2.2.  Integrated path stability selection.
-In Section 2.1, we showed that efp scores can be used to control E(FP), the FDR, and $q$-values.
-In this section, we review integrated path stability selection (IPSS), which constructs efp scores from arbitrary feature selection algorithms.
-The advantage of IPSS over other stability selection methods is that its efp scores have the tightest E(FP) bounds (Melikechi and Miller, 2024).
-Let $S$ be the unknown subset of important features as before, and let $\hat{S}_{\lambda}\subseteq\{1,\ldots,p\}$ be an estimator of $S$ that depends on the data and a parameter $\lambda>0$, with larger values of $\lambda$ corresponding to fewer features being selected.
-Importantly, $\hat{S}_{\lambda}$ and $\hat{S}(t)$ are distinct estimators of $S$: The former is a baseline selection algorithm whose parameter $\lambda$ will always appear as a subscript, and the latter is an estimator based on efp scores whose parameter $t$ will always appear as a functional argument.
-Stability selection uses subsampling to construct $\hat{S}(t)$ from $\hat{S}_{\lambda}$ (Meinshausen and Bühlmann, 2010; Shah and Samworth, 2013).
-Specifically, rather than estimate $S$ using all of the data at once, one randomly draws disjoint subsets $A_{1},A_{2}\subseteq\{1,\dots,n\}$ of size $\lfloor n/2\rfloor$ and evaluates $\hat{S}_{\lambda}(Z_{A_{1}})$ and $\hat{S}_{\lambda}(Z_{A_{2}})$ at all $\lambda$ in some interval $\Lambda\subseteq(0,\infty)$, where $Z_{A}=(Z_{i}:i\in A)$.
-After $B$ subsampling iterations, the estimated selection probability $\hat{\pi}_{j}(\lambda)$ is the proportion of times feature $j$ is selected over all $2B$ subsets, $\hat{\pi}_{j}(\lambda)=\frac{1}{2B}\sum_{b=1}^{2B}\operatorname{\mathds{1}}(j\in\hat{S}_{\lambda}(Z_{A_{b}}))$.
+2.2  Integrated path stability selection.
+Integrated path stability selection (IPSS) constructs efp scores by applying baseline feature selection algorithms to random subsamples of the data.
+Specifically, let $S$ be an unknown subset of true features as before, and let $\hat{S}_{\lambda}\subseteq\{1,\ldots,p\}$ be an estimator of $S$ that depends on the data and a parameter $\lambda>0$.
+Note that $\hat{S}_{\lambda}$ and $\hat{S}(t)$ are distinct estimators of $S$: The former is a baseline algorithm whose parameter $\lambda$ appears as a subscript.
+The IPSS subsampling procedure, which is also used in stability selection, consists of $B$ subsampling iterations.
+Each iteration consists of randomly drawing disjoint subsets $A_{1},A_{2}\subseteq\{1,\dots,n\}$ of size $\lfloor n/2\rfloor$ and evaluating $\hat{S}_{\lambda}(Z_{A_{1}})$ and $\hat{S}_{\lambda}(Z_{A_{2}})$ at all $\lambda$ in some interval $\Lambda\subseteq(0,\infty)$, where $Z_{A}=(Z_{i}:i\in A)$ (the choice of $\lfloor n/2\rfloor$ samples is needed for existing stability selection theorems—not just IPSS—to hold).
+After $B$ iterations, the estimated selection probability $\hat{\pi}_{j}(\lambda)=\frac{1}{2B}\sum_{b=1}^{2B}\operatorname{\mathds{1}}(j\in\hat{S}_{\lambda}(Z_{A_{b}}))$ of feature $j$ is the proportion of times $j$ is selected over all $2B$ subsets.
+Large values of $\hat{\pi}_{j}(\lambda)$ correspond to $j$ being selected by $\hat{S}_{\lambda}$ on many of the random subsamples, suggesting that $j$ is important.
 Melikechi and Miller (2024) prove that for any $\Lambda\subseteq(0,\infty)$, any probability measure $\mu$ on $\Lambda$, and certain functions $f:[0,1]\to\mathbb{R}$, the function $\mathtt{efp}_{Z_{1:n}}:\{1,\dots,p\}\to[0,p]$ defined by
 $\mathtt{efp}_{Z_{1:n}}(j)$ $=\min\left\{\frac{\mathcal{I}(\Lambda)}{\int_{\Lambda}f(\hat{\pi}_{j}(\lambda))\mu(d\lambda)},\ p\right\},$ (2.3)
-is a valid efp score, where $\mathcal{I}(\Lambda)$ is a parameter that is described below and in Section S1.
-While several choices of $f$ are possible in Equation 2.3, we always use $f(x)=(2x-1)^{3}\operatorname{\mathds{1}}(x\geq 0.5)$ because the resulting efp score bounds $\mathrm{E}(\mathrm{FP}(t))\leq t$ are the tightest of any existing version of stability selection (Melikechi and Miller, 2024).
-[…]
-2.3.  IPSS for thresholding.
+is a valid efp score, that is, $\hat{S}(t)=\{j:\mathtt{efp}_{Z_{1:n}}(j)\leq t\}$ satisfies $\mathrm{E}\lvert\hat{S}(t)\cap S^{c}\rvert\leq t$.
+An explicit form of $\mathcal{I}(\Lambda)$, which comes from the theory of IPSS, is provided in Section S1.
+Several choices of $f$ in Equation 2.3 yield provably valid efp scores (see Section S1 as well as Theorems 4.1, 4.2, S3.2, S3.3 in Melikechi and Miller (2024) for detailed theoretical results about IPSS with different functions).
+We always use $f(x)=(2x-1)^{3}\operatorname{\mathds{1}}(x\geq 0.5)$ because the resulting bounds $\mathrm{E}(\mathrm{FP}(t))\leq t$ are the tightest of any existing version of stability selection (Melikechi and Miller, 2024).
+In addition to its theoretical justification, the empirical results in Figures S18 and S19 show that this choice of $f$ produces the best results among several functions for which valid efp scores are available.
+Further details about IPSS, including descriptions of $\Lambda$, $\mu$, and a derivation of Equation 2.3, are provided in Section S1.
+2.3  IPSS for feature importance scores.
+Until now, IPSS has only been applied to parametric baseline estimators $\hat{S}_{\lambda}$.
+A canonical example is $\ell^{1}$-regularized estimators $\hat{S}_{\lambda}=\{j:\hat{\beta}_{j}(\lambda)\neq 0\}$, where
+$\hat{\beta}(\lambda)$ $=\operatorname*{arg\,min}_{\beta\in\mathbb{R}^{p}}\sum_{i=1}^{n}\mathcal{L}(Y_{i},\beta^{\mathtt{T}}X_{i})+\lambda\sum_{j=1}^{p}\lvert\beta_{j}\rvert.$ (2.4)
+Here, $\lambda>0$ controls the strength of the $\ell^{1}$-penalty and $\mathcal{L}$ is a log-likelihood that assumes a specific relationship between the features and response.
+By extending IPSS to feature importance scores, we no longer require such restrictive assumptions.
 An importance function is a (possibly random) map $\Phi_{Z_{1:n}}:\{1,\dots,p\}\to\mathbb{R}$ that uses the data to assign an importance score $\Phi_{Z_{1:n}}(j)$ to each feature.
 The possible randomness, which is additional to the randomness in $Z_{1:n}$, can come from, for example, random subsampling in tree-based algorithms.
-Suppressing $Z_{1:n}$ from the notation for now, assume that $\Phi(j)<\Phi(k)$ means $j$ is less important than $k$ according to $\Phi$.
-For example, in linear regression where $Z_{i}=(X_{i},Y_{i})$ with $Y_{i}=\beta^{\mathtt{T}}X_{i}+\epsilon_{i}$ and $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$, a viable importance function is the magnitude of the estimated regression coefficient, $\Phi(j)=\lvert\hat{\beta}_{j}\rvert$.
-Another class of importance functions—the focus in this work—come from tree ensemble methods such as gradient boosted trees and random forests, where again we are in the supervised setting $Z_{i}=(X_{i},Y_{i})$ (Friedman, 2001; Breiman, 2001).
-Given a collection of binary decision trees, $\mathcal{T}$, these importance functions take the form
-$\Phi(j)$ $=\frac{1}{\lvert\mathcal{T}\rvert}\sum_{T\in\mathcal{T}}\sum_{v\in T}\Delta\varphi(v)\operatorname{\mathds{1}}(j=j_{v})$ (2.4)
-where the outer sum is over all trees $T\in\mathcal{T}$, the inner sum is over all nodes $v\in T$, $j_{v}$ is the feature used to split node $v$, and $\varphi(v)$ measures the impurity of $v$.
-For regression, we use the squared error loss, $\varphi(v)=\sum_{i\in v}(Y_{i}-\overline{Y}_{v})^{2}/\lvert v\rvert$, where $v\subseteq\{1,\dots,n\}$ is identified with the subset of samples in node $v$, and $\overline{Y}_{v}=\sum_{i\in v}Y_{i}/\lvert v\rvert$ is the empirical mean of the responses in $v$.
-For binary classification, we use the Gini index, $\varphi(v)=2p_{0}(v)p_{1}(v)$ where, for $a\in\{0,1\}$, $p_{a}(v)=\sum_{i\in v}\operatorname{\mathds{1}}(Y_{i}=a)/\lvert v\rvert$ is the proportion of responses in $v$ that equal $a$.
-The change in impurity
-$\Delta\varphi(v)$ $=\varphi(v)-\bigg{(}\frac{\lvert v_{L}\rvert}{\lvert v\rvert}\varphi(v_{L})+\frac{\lvert v_{R}\rvert}{\lvert v\rvert}\varphi(v_{R})\bigg{)}$
-is the impurity difference between $v$ and its children, $v_{L}$ and $v_{R}$.
-Large positive values of $\Delta\varphi(v)$ indicate that the feature $j_{v}$ used to split node $v$ successfully partitions the data in a manner that is consistent with the objective of the tree.
-Importance scores naturally lend themselves to thresholding: Given a $\Phi$ and a threshold $\lambda$, the selected set of features is $\hat{S}_{\lambda}=\{j:\Phi(j)\geq\lambda\}$.
-Algorithm S1 outlines IPSS for thresholding.
+Suppressing $Z_{1:n}$ from the notation for now, assume $\Phi(j)<\Phi(k)$ means that $j$ is less important than $k$ according to $\Phi$.
+For example, in linear regression where $Z_{i}=(X_{i},Y_{i})$ with $Y_{i}=\beta^{\mathtt{T}}X_{i}+\epsilon_{i}$, a viable importance function is the magnitude of the estimated regression coefficient, $\Phi(j)=\lvert\hat{\beta}_{j}\rvert$.
+Associated to any importance function $\Phi$ is a baseline feature selection estimator $\hat{S}_{\lambda}=\{j:\Phi(j)\geq\lambda\}$, obtained by simply thresholding the importance scores.
+That is, given $\Phi$ and a threshold $\lambda$, $\hat{S}_{\lambda}$ selects the features whose importance scores are at least $\lambda$.
+With $\hat{S}_{\lambda}$ defined, we have all that is needed to implement IPSS.
+Furthermore, all of the theoretical results from Melikechi and Miller (2024) apply without modification.
+2.3.1  Implementation.
+Algorithm S1 outlines IPSS for feature importances.
 The main steps are as follows: First, features are preselected according to the procedure described in Section S2.1.
@@
 The largest, $\lambda_{\mathrm{max}}$, is the maximum importance score across all features and all $2B$ sets of scores (hence, $\hat{S}_{\lambda_{\mathrm{max}}}=\varnothing$).
-Next, starting from $\lambda_{\mathrm{min}}=\lambda_{\mathrm{max}}$, we decrease $\lambda_{\mathrm{min}}$ one step at a time, usually on a log scale, iteratively updating the integral $\mathcal{I}([\lambda_{\mathrm{min}},\lambda_{\mathrm{max}}])$ at each step in the form of a Riemann sum approximation until $\mathcal{I}([\lambda_{\mathrm{min}},\lambda_{\mathrm{max}}])$ surpasses the cutoff $C=0.05$.
+Next, starting from $\lambda_{\mathrm{min}}=\lambda_{\mathrm{max}}$, decrease $\lambda_{\mathrm{min}}$ one step at a time, usually on a log scale, iteratively updating the integral $\mathcal{I}([\lambda_{\mathrm{min}},\lambda_{\mathrm{max}}])$ at each step in the form of a Riemann sum approximation until $\mathcal{I}([\lambda_{\mathrm{min}},\lambda_{\mathrm{max}}])$ surpasses a cutoff $C$.
 Once $C$ is surpassed, the $\mathtt{while}$ loop stops and the feature-specific selection probabilities and integrals are evaluated.
-The algorithm outputs the efp scores for each feature.
-Given a target E(FP), $t$, the set of selected features is $\hat{S}(t)=\{j:\mathtt{efp}_{Z_{1:n}}(j)\leq t\}$.
-2.4.  IPSSGB and IPSSRF.
-We consider two instances of the general method: IPSSGB uses gradient boosting for the baseline algorithm $\hat{S}_{\lambda}$, and IPSSRF uses random forests for $\hat{S}_{\lambda}$.
-All other IPSS parameters are identical, except we use $B=50$ for IPSSRF and $B=100$ for IPSSGB.
-The latter led to slightly more stable results with hardly any increase in runtimes (no such improvement was observed for IPSSRF).
-In all that follows, random forests are implemented with scikit-learn (Pedregosa et al., 2011) and gradient boosting with XGBoost (Chen and Guestrin, 2016).
-For random forests, all parameters are set to their default settings except for the proportion of features considered when splitting each node (called $\mathtt{max\_features}$ in scikit-learn and often $\mathtt{mtry}$ elsewhere), which we change from $1$ to $1/3$.
-This is a standard choice (Biau and Scornet, 2016) and led to an approximately $3$-fold speedup in IPSSRF without any discernible effect on the results.
-For gradient boosting, we also change the proportion of features considered at each split from $1$ to $1/3$.
-Aside from this, the only non-default parameter is the maximum depth of each tree, which we change from $6$ to $1$, making each tree a stump.
-This significantly improved the performance of IPSSGB, both in terms of speed and feature selection.
-Finally, while many different importance functions exist for gradient boosting and random forests, we use mean decrease impurity, Equation 2.4, with the squared error loss for regression and the Gini index for classification, as defined in Section 2.3.
-Experimentation with other importance functions showed similar results, though a more detailed study of IPSS with different importance functions is a potentially direction for future work.
-2.4.1.  Further background.
-The seminal work on gradient boosting is Friedman (2001).
-For random forests, we refer to the original work of Breiman (2001), and to Biau and Scornet (2016) for an excellent review with a detailed description of importance scores.
-Gradient boosting, random forests, and their importance scores are also detailed in Hastie et al.
-(2009).
-Finally, our formulation of tree-based scores partly follows Louppe et al.
-(2013), who provide a clear and concise formulation of these quantities.
-3.  Other methods
-We describe several feature selection methods that are compared to IPSSGB and IPSSRF in Sections 4 and 5.
-We selected these methods due to their popularity and superior performance relative to other methods in several comparison studies (Degenhardt et al., 2019; Sanchez-Pinto et al., 2018; Speiser et al., 2019).
-3.1.  Methods with error control.
-The closest method to IPSSGB in terms of its underlying approach is that of Hofner et al.
-(2015), referred to here as SSBoost.
-Unlike IPSSGB—which uses importance scores from gradient boosting—SSBoost applies stability selection to choose the number of features used per boosting run.
-Furthermore, IPSSGB uses IPSS to construct efp scores, whereas SSBoost uses a version of stability selection introduced by Shah and Samworth (2013).
-This is perhaps the most significant difference since the efp scores for IPSS have much tighter bounds than those for other forms of stability selection (Melikechi and Miller, 2024).
-From a practical standpoint, this causes other versions of stability selection to identify fewer important features than IPSS for the same target E(FP).
-IPSS has previously been used by Melikechi and Miller (2024) with lasso for regression and $\ell^{1}$-regularized logistic regression for classification (Tibshirani, 1996; Lee et al., 2006).
-These methods, abbreviated as IPSSL, are part of a more general class of $\ell^{1}$-regularized estimation techniques to which stability selection is commonly applied (Meinshausen and Bühlmann, 2010), namely,
-$\hat{\beta}$ $=\operatorname*{arg\,min}_{\beta\in\mathbb{R}^{p}}\sum_{i=1}^{n}\mathcal{L}(Y_{i},\beta^{\mathtt{T}}X_{i})+\lambda\sum_{j=1}^{p}\lvert\beta_{j}\rvert,$
-where $\mathcal{L}$ is a log-likelihood function and $\lambda>0$ controls the strength of the $\ell^{1}$ penalty.
-By contrast, since decision trees are nonparametric, neither IPSSGB nor IPSSRF assume an underlying model.
-Furthermore, unlike linear regression and logistic regression, decision trees can effectively capture nonlinear relationships in data.
-The nonlinear and nonparametric nature of decision trees motivated our extension of IPSS to gradient boosting and random forests.
-In Section 4, we find that both methods significantly outperform IPSSL in the presence of nonlinearity.
-Another method to which we compare is knockoff boosted trees (KOBT) (Jiang et al., 2021).
-KOBT applies the model-X knockoff framework of Candes et al.
-(2018) to boosted trees to obtain a method with theoretical FDR control.
-We use the R package KOBT, preselecting $\mathtt{bound}=200$ features and generating $\mathtt{num}=100$ knockoff matrices.
-The knockoff type is shrunken Gaussian.
-These choices were made to balance performance and runtime.
-Computation times were much longer for larger values of $\mathtt{bound}$ or $\mathtt{num}$, or when using a different version of KOBT that employs principal components to avoid the Gaussian assumption.
-Moreover, we found that these computationally more expensive choices did not noticeably improve performance or FDR control.
-We also tested the method of Coleman et al.
-(2022), which achieves theoretical error control by using random forests for hypothesis testing.
-However, one test run with default parameters on simulated data with 500 samples, 500 features, and 20 true features took over 51 minutes, returning 15 true positives and 43 false positives.
-By contrast, IPSSGB with a target E(FP) of 2 took 11 seconds and returned 10 true positives and 0 false positives on the same data.
-This method is omitted from our study because its performance does not appear to justify its excessive runtime.
-3.2.  Methods without error control.
-We also compare to several popular methods without theoretical error control: Boruta (Kursa et al., 2010), random forests (RF) (Breiman, 2001), Vita (Janitza et al., 2018), VSURF (Genuer et al., 2010), and XGBoost (Chen and Guestrin, 2016).
-Boruta, Vita, and VSURF use random forests.
-In terms of performance, Speiser et al.
-(2019) found Boruta and VSURF were among the best out of $13$ random forest-based methods, and Degenhardt et al.
-(2019) found Boruta and Vita outperformed $5$ other methods on several metrics.
-Method Package Error control Base method Non-default settings
-IPSSGB ipss (Python) Boosting
-IPSSRF ipss (Python) Random forest
-IPSSL ipss (Python) Lasso
-SSBoost XGBoost (Python) Boosting $\tau=0.75$
-with stabs (R)
-KOBT KOBT (R) Boosting $\texttt{num}=100$, $\texttt{bound}=200$,
+The algorithm outputs efp scores for each feature, which are used to control E(FP), FDR, and compute $q$-values, as described in Section 2.1.
 […]
-Boruta Boruta (R) Random forest
-RF scikit-learn (Python) Random forest
-Vita vita (R) Random forest $p$-value threshold $=0$
-VSURF VSURF (R) Random forest VSURF_pred
-XGBoost XGBoost (Python) Boosting
-Table 1.
-Feature selection methods.
-Packages, when available, are listed with the language used to implement them in parentheses.
-Non-default settings are elaborated upon in Sections 3.1 and 3.2.
-For methods with no non-default settings, we use the default settings in their respective packages.
-Table 1 summarizes the methods.
-Details regarding the implementation and choice of settings are as follows.
-Hofner et al.
-(2015) provide code for SSBoost that combines the R packages mboost (Hofner et al., 2014) and stabs (Hofner and Hothorn, 2017) in the form of a worked example, but the mboost implementation of boosting was prohibitively slow in the dimensions we consider.
-By adapting their code to use XGBoost in place of mboost for boosting and by preselecting features exactly as we do for IPSSGB (Section S1.4), we were able to reduce SSBoost runtimes considerably with no apparent change in results.
-The XGBoost parameters used for SSBoost are the same as those used for IPSSGB; specifically, all XGBoost parameters are set to their default values except the proportion of features considered at each split (changed from 1 to 1/3) and the depth of each tree (changed from 6 to 1).
-For the stability selection part of SSBoost we use the default parameters in stabs, and the selection threshold is set to $\tau=0.75$, which is the middle of the interval $(0.6,0.9)$ recommended by Meinshausen and Bühlmann (2010).
-For Vita, we set the $p$-value threshold to 00, and for VSURF we use the function VSURF_pred rather than VSRUF_interp to select the final set of features (Genuer et al., 2010); both choices correspond to as few features as possible being selected under the constraints of the respective algorithms, a favorable quality due to the sparsity in our forthcoming studies.
-For XGBoost and RF, we run the respective algorithms on the full dataset with default parameters, then select the top ranked features by their importance scores.
-4.  Simulation study
-In this section, we conduct a simulation study to evaluate the performance of the methods in Table 1 when the true set of important features is known.
-To make the simulations more realistic, we use features from real data rather than generating them from known distributions.
-Specifically, given $n$ and $p$, in each trial we randomly select $n$ samples and $p$ features from RNA sequencing (RNA-seq) data from 569 ovarian cancer patients (samples) and 6426 genes (features) (Vasaikar et al., 2018).
-This publicly available dataset, part of The Cancer Genome Atlas (Weinstein et al., 2013), was chosen because it is high dimensional and the features follow a variety of distributions.
-For instance, Figure 2 shows the standardized empirical distributions of five randomly selected genes, which, from left to right, are relatively flat, skewed left, approximately Gaussian, skewed right, and contain outliers.
+2.3.2  IPSSGB and IPSSRF.
+Any importance function can be combined with IPSS, providing many possible directions for future work.
+In this paper, we focus on importance functions from tree ensemble methods because they are nonparametric, computationally efficient, and produce state-of-the-art results on tabular data (Shwartz-Ziv and Armon, 2022).
+These importance functions are defined as follows; for additional details, see Louppe et al.
+(2013) or Biau and Scornet (2016).
+Given a collection of binary decision trees, $\mathcal{T}$, define
+$\Phi(j)$ $=\frac{1}{\lvert\mathcal{T}\rvert}\sum_{T\in\mathcal{T}}\sum_{v\in T}\Delta\varphi(v)\operatorname{\mathds{1}}(j=j_{v})$ (2.5)
+where the outer sum is over all trees $T\in\mathcal{T}$, the inner sum is over all nodes $v\in T$, $j_{v}$ is the feature used to split node $v$, and $\varphi(v)$ measures the impurity of $v$.
+The change in impurity
+$\Delta\varphi(v)$ $=\varphi(v)-\bigg{(}\frac{\lvert v_{L}\rvert}{\lvert v\rvert}\varphi(v_{L})+\frac{\lvert v_{R}\rvert}{\lvert v\rvert}\varphi(v_{R})\bigg{)}$
+is the impurity difference between $v$ and its children, $v_{L}$ and $v_{R}$.
+Large positive values of $\Delta\varphi(v)$ indicate that the feature $j_{v}$ used to split node $v$ successfully partitions the data in a manner that is consistent with the objective of the tree.
+For regression, we use the squared error loss impurity function, $\varphi(v)=\sum_{i\in v}(Y_{i}-\overline{Y}_{v})^{2}/\lvert v\rvert$, where $v\subseteq\{1,\dots,n\}$ is identified with the subset of samples in node $v$, and $\overline{Y}_{v}=\sum_{i\in v}Y_{i}/\lvert v\rvert$ is the empirical mean of the responses in $v$.
+For binary classification, we use the Gini index, $\varphi(v)=2p_{0}(v)p_{1}(v)$ where, for $a\in\{0,1\}$, $p_{a}(v)=\sum_{i\in v}\operatorname{\mathds{1}}(Y_{i}=a)/\lvert v\rvert$ is the proportion of responses in $v$ that equal $a$.
+These were selected because they are canonical choices and because we found little difference in results between these and other standard impurity functions.
+The importance function defined in Equation 2.5 can be computed for any tree ensemble method.
+We focus on gradient boosting (Friedman, 2001) and random forests (Breiman, 2001).
+As noted above, these are abbreviated IPSSGB and IPSSRF when combined with IPSS.
+Additional implementation details about IPSSGB and IPSSRF are in Sections S2.2.1 and S2.2.2, respectively.
+2.3.3  Other importance functions.
+The importance function in Equation 2.5 is called mean decrease impurity (MDI) because it measures the average decrease in impurity attributed to each feature over all trees.
+Another common importance function is mean decrease accuracy (MDA), also known as permutation importance (Louppe et al., 2013).
+MDA is more general than MDI in that it can be applied to any predictive model, not just tree ensembles.
+Several variants of MDA exist, but the basic procedure is: (i) train the model, (ii) compute the prediction error $e$ on test data, and (iii) for each feature $j$, randomly permute the values of feature $j$ in the test data (keeping all other features unchanged), and compute the prediction error $e_{j}$ of the trained model on the permuted test data.
+The underlying idea, captured by the importance score $\Phi(j)=e_{j}-e$, is that permuting unimportant features should have little effect on the prediction error, while permuting important features should degrade predictive performance.
+We experimented with IPSS applied to both MDI and MDA from gradient boosting and random forests.
+In both cases, the FDR was controlled at target levels, but IPSS with MDI consistently identified more true positives.
+One likely reason for this is that MDA tends to spread importances more uniformly across features than MDI, making it more difficult to distinguish between important and unimportant features (Hastie et al., 2009).
+Furthermore, MDI, whose importance scores are obtained during the training process itself, is more computationally efficient than MDA, which requires the additional step of evaluating the trained model on the permuted data for every feature.
+The recent success of deep learning, especially on text and image data, has generated much interest in feature importance scores—for example, SHAP values (Lundberg and Lee, 2017)—that are derived from these models.
+Applying IPSS to such scores is a potentially interesting line of future work.
+However, numerous studies have shown that tree ensemble methods like XGBoost (Chen and Guestrin, 2016) are faster, easier to train, and consistently outperform deep learning methods on tabular data, especially when there are fewer than 10,000 samples (Shwartz-Ziv and Armon, 2022; Grinsztajn et al., 2022; Fayaz et al., 2022).
+Since our focus in this work is on tabular medical data, where the number of patients rarely exceeds several hundred, we primarily consider importance scores from tree ensemble methods given their many advantages over deep learning in this setting.
+2.3.4  Computation.
+The IPSS subsampling procedure requires $2B$ feature importance evaluations, one evaluation on each half of the data in all $B$ iterations.
+This is fast for MDI scores; when combined with preselection (Section S2.1), IPSSGB and IPSSRF run in under 20 seconds on a standard laptop when there are $n=500$ samples and $p=5000$ features (Tables S3 and S4).
+Since this is already sufficiently fast for our purposes, we do not implement more efficient alternatives in this work.
+We note, however, that IPSS is embarrassingly parallel: All iterations of the subsampling procedure can be evaluated separately and hence run simultaneously, potentially accelerating IPSS by a factor of $2B$ (typically 100 or 200).
+This is especially beneficial when importance scores are expensive to compute, as is often the case when they come from deep learning methods.
+By contrast, many feature selection methods are iterative and thus not parallelizable.
+For example, recursive feature elimination (Li et al., 2022) typically removes features one at a time based on predictive performance, requiring the model to be retrained after each removal.
+When $p$ is large, this can require thousands of model fits, far more than the $2B$ evaluations typically required by IPSS.
+3.  Simulation studies
+In this section, we conduct two simulation studies to evaluate the performance of 14 feature selection methods when the true set of important features is known.
+Features in the first study are drawn from a multivariate Gaussian, and the response is generated from a nonlinear additive model.
+To make the simulations more realistic, in the second study we use features from real RNA sequencing (RNA-seq) data rather than generating them from known distributions, and the response is a highly randomized, nonlinear function of the important features.
+3.1  Other methods.
+We compare IPSSGB and IPSSRF to 12 feature selection methods: IPSS with $\ell^{1}$-regularization (IPSSL1, Melikechi and Miller (2024)); boosting with stability selection (SSBoost, Hofner et al.
+(2015)); five versions of model-X knockoffs (Candes et al., 2018), namely knockoffs with generalized linear models (KOGLM), knockoffs with lasso (KOL1), knockoffs with random forests (KORF), knockoffs with boosted trees (KOBT, Jiang et al.
+(2021)), and knockoffs with deep neural networks (DeepPINK, Lu et al.
+(2018)); random forest hypothesis testing (RFHT, Coleman et al.
+(2022)); Boruta (Kursa et al., 2010); recursive feature elimination with gradient boosting (RFEGB); Vita (Janitza et al., 2018); and VSURF (Genuer et al., 2010).
+We provide a brief overview of these methods below.
+Additional details, including parameter settings and the software packages used to implement each method, are in Section S2.
+Four methods—Boruta, RFEGB, Vita, and VSURF—do not have theoretical false discovery control.
+We chose these because Speiser et al.
+(2019) found that Boruta and VSURF were among the best out of $13$ random forest-based feature selection methods, and Degenhardt et al.
+(2019) found that Boruta and Vita outperformed $5$ other methods in extensive comparison studies.
+We tested RFEGB so as to include a gradient boosting-based feature selection method without false discovery control.
+IPSSL1 and SSBoost provide theoretical E(FP) control.
+IPSSL1, a parametric version of IPSS, assumes a (generalized) linear relationship between the features and the response (Equation 2.4).
+SSBoost combines gradient boosting and stability selection.
+We implement it assuming the $r$-concavity conditions of Shah and Samworth (2013), which are required to obtain the tightest upper bound on E(FP) of any version of stability selection other than IPSS (note that IPSS does not require $r$-concavity assumptions) (Melikechi and Miller, 2024).
+Model-X knockoffs is a general framework for feature selection with theoretical FDR control that has attracted much attention recently, in part due to its flexibility.
+Like IPSS, model-X knockoffs works in high-dimensions ($p>n$), does not use $p$-values (which are challenging to compute in general), is compatible with arbitrary feature importance scores, and makes no assumptions about the relationship between the response and the features.
+Unlike IPSS, model-X knockoffs assumes that the joint distribution of the features is known (Candes et al., 2018).
+3.2  Gaussian simulation design.
+We perform two regression experiments and two classification experiments.
+The two experiments in both settings correspond to $n=250$ and $500$.
+All experiments consist of $100$ trials.
+In each trial, $n$ independent samples of $X\sim\mathcal{N}(0,\Sigma)$ are drawn from a $p=500$-dimensional, mean-zero multivariate Gaussian with a Toeplitz covariance matrix, $\Sigma$.
+The correlation parameter of the Toeplitz matrix is set to $\rho=0.5$; that is, $\Sigma_{jk}=0.5^{\lvert j-k\rvert}$.
+The number of important features $\lvert S\rvert$ is drawn uniformly at random from $\{5,\dots,15\}$ prior to each trial, and a new set of important features $S$ of size $\lvert S\rvert$ is randomly selected.
+In each experiment, the signal is $f(X)=\sum_{j\in S}\exp(-X_{j}^{2})$.
+For regression, the response is $Y=f(X)+\epsilon$ where $\epsilon\sim\mathcal{N}(0,\sigma^{2})$ and $\sigma^{2}$ is selected according to a specified signal-to-noise ratio (SNR) that is drawn uniformly at random from the interval $[0.5,2]$.
+For classification, we draw $Y\sim\mathrm{Bernoulli}(\pi)$ where $\pi=1/(1+\exp(-uf(X)))$ and the signal strength $u$ is drawn uniformly at random from the interval $[1,3]$.
+These sources of randomness are introduced so that the study covers a wide range of settings.
+Figure 1.
+Gaussian simulation results ($n=500$).
+First and second columns show the regression and classification results, respectively.
+The horizontal axis in each plot shows the target FDR.
+The three methods without false discovery control (gray horizontal lines) do not vary with the target FDR.
+The black dashed line in the FDR plots represents perfect FDR calibration, FDR = target FDR.
+Non-gray solid lines correspond to IPSS or stability selection-based methods.
+Non-gray dashed lines show methods based on model-X knockoffs.
+3.3  RNA-seq simulation design.
+We perform three regression experiments and three classification experiments.
+The three experiments in both settings correspond to $p=500$, $2000$, and $5000$.
+All experiments consist of $100$ trials.
+In each trial, $n=500$ patients and $p$ genes are randomly selected from RNA-seq measurements of 6426 genes from $569$ ovarian cancer patients (Vasaikar et al., 2018).
+This publicly available dataset, part of The Cancer Genome Atlas (Weinstein et al., 2013), was chosen because it is high dimensional and the features follow a variety of empirical distributions (Figure S5).
 Furthermore, the genes exhibit complex correlation structures, with maximum and average absolute pairwise correlations of approximately $0.95$ and $0.17$ after standardization, respectively.
-[…]
-Figure 1.
-Simulating the response.
-Random rows and columns are selected from the full RNA-seq dataset to create an $n$-by-$p$ design matrix $X$, whose columns are then split into important features, $X_{S}$, and unimportant features, $X_{S^{c}}$.
-The columns of $X_{S}$ are further partitioned into $G$ groups; the above figure shows the case $G=3$.
-The features in each group are summed to obtain $\xi_{g}$, and a different realization $f_{\theta_{g}}$ of $f_{\theta}$ is applied to $\xi_{g}$ for each $g$.
-The response is the sum of the resulting group-specific signals, plus noise.
-Figure 2. Distributions of five genes from the RNA-seq dataset.
-4.1.  Simulation details.
-Algorithm S2 describes the simulation procedure, which is also illustrated in Figure 1.
+Algorithm S2 describes the simulation procedure for each individual trial, which is also illustrated in Figure S3.
 In all steps, “randomly select” means select a parameter uniformly at random from its domain.
 The general outline is as follows: First, randomly select an $n$-by-$p$ submatrix $X$ of the full RNA-seq dataset and standardize its columns to have mean 0 and variance 1.
-Next, randomly select a true subset of $p_{\mathrm{true}}$ important features $S$, and partition it into $G$ groups, $S_{1},\dots,S_{G}$.
+Next, the number of important features $\lvert S\rvert$ is drawn uniformly at random from $\{10,\dots,30\}$ and a randomly selected subset of $\lvert S\rvert$ important features $S$ is partitioned into $G$ groups, $S_{1},\dots,S_{G}$.
 A different realization of a randomized nonlinear function $f_{\theta_{g}}$, defined in Equation S3.1, is applied to the standardized sum $\xi_{g}=(\hat{\xi}_{g}-\bar{\xi}_{g})/\bar{\sigma}_{g}$ of the features in each group $S_{g}$, where $\bar{\xi}_{g}$ and $\bar{\sigma}_{g}$ are the empirical mean and standard deviation of $\hat{\xi}_{g}=\sum_{j\in S_{g}}X_{j}$.
@@
 This scheme produces data with highly complex interactions between features and the response, going well beyond the additive setting $Y=\sum_{j\in S}f_{j}(X_{j})+\epsilon$.
-Regression and classification data are simulated with $p=500$, $2000$, and $5000$ features for a total of six experiments.
-Each experiment consists of $100$ trials, outlined in Algorithm S2.
-Each trial uses $n=500$ samples and $\lvert S\rvert=p_{\mathrm{true}}$ important features.
-For regression, the response is $y=\sum_{g=1}^{G}f_{\theta}(\xi_{g})+\epsilon$, where $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$ and the variance $\sigma^{2}$ is selected according to a specified signal-to-noise ratio (SNR).
-For classification, we draw $y_{i}\sim\mathrm{Bernoulli}(\pi_{i})$ where $\pi_{i}=1/(1+\exp(-u\eta_{i}))$.
-A new $p_{\mathrm{true}}$ and SNR (or $u$), as well as new sets of samples, features, and important features, are drawn prior to each trial, with $p_{\mathrm{true}}$, SNR, and $u$ chosen according to Table S1.
-The many sources of randomness in our setup are introduced so that the simulations cover a wide range of settings.
-4.2.  Simulation results.
-Figures 3, LABEL:, 4, LABEL:, and 5 show the results.
-The empirical false discovery and true positive rates are $\mathrm{FDR}=\mathrm{FP}/(\mathrm{TP}+\mathrm{FP})$ and $\mathrm{TPR}=\mathrm{TP}/(\mathrm{TP}+\mathrm{FN})$, where FP, TP, and FN are the number of false positives, true positives, and false negatives, respectively.
-The upper left (high TP or TPR and low FP or FDR) corresponds to better performance in Figures 3 and 4.
-The ability of the different methods to control E(FP) (Figure 3) and FDR (Figure 4) at their target values is shown when applicable.
-Specifically, SSBoost and KOBT provide theoretical control of E(FP) and FDR, respectively, while the IPSS methods can be specified to control either E(FP) or FDR (Section 2.2).
-In Figure 3, the dashed vertical line signifies a target E(FP) of 3, with symbols showing the average FP and TP for the relevant methods at this target value.
-In particular, the for a method with perfect E(FP) control should lie on the dashed vertical line.
-Similarly, the and symbols in Figure 4 correspond to target FDRs of 0.1 and 0.3, respectively.
+For regression, the response is $Y=\eta+\epsilon$, where $\epsilon\sim\mathcal{N}(0,\sigma^{2})$ and the variance $\sigma^{2}$ is selected according to a specified signal-to-noise ratio (SNR) that is drawn uniformly at random from the interval $[0.5,2]$.
+For classification, we draw $Y\sim\mathrm{Bernoulli}(\pi)$ where $\pi=1/(1+\exp(-u\eta))$ and the signal strength $u$ is drawn uniformly at random from $[1,3]$.
+As in the Gaussian simulation experiments, these many sources of randomness are introduced to cover a wide range of settings.
+3.4  Simulation results.
+Figures 1 and S1 show the results of the Gaussian simulations when $n=500$ and $250$, respectively, and Figure 2 and S2 show the results of the RNA-seq simulations for regression and classification.
+Runtimes for each method in each experiment are provided in Tables S3 and S4.
+The FDR in each plot is the average of $\mathrm{FP}/(\mathrm{TP}+\mathrm{FP})$ over all 100 trials, and the true positive rate (TPR) is the average of $\mathrm{TP}/(\mathrm{TP}+\mathrm{FN})$, where FP, TP, and FN are the number of false positives, true positives, and false negatives, respectively.
+Both FDR and TPR are shown as functions of the target FDR.
+The black dashed line in each FDR plot represents perfect FDR calibration, $\mathrm{FDR}=\mathrm{target\ FDR}$.
+A method’s FDR is well-controlled if its FDR lies on or below this line.
+The FDR and TPR for methods without false discovery control are shown as horizontal lines because they do not admit false discovery control parameters and therefore do not vary with the target FDR.
 IPSSGB has the best performance overall.
-Its curves are the farthest to the upper left in every plot in Figures 3 and 4, and it has the best control of both E(FP) and FDR.
-In Figure 3, for example, the for IPSSGB is the closest to the vertical dashed line in all experiments, and similarly in Figure 4.
-Furthermore, IPSSGB identifies the most true positives, both in terms of average TP and average TPR.
-Notably, at a target E(FP) of 3 and a target FDR of 0.3, IPSSGB has the largest average TP and highest TPR in nearly every setting, while having significantly fewer false positives than most other methods.
-Figure 5 also shows that IPSSGB is one of the fastest methods, running in under 20 seconds in all settings.
+Its FDR is always well-controlled and it consistently has a much higher TPR than the other methods with false discovery control.
+Notably, IPSSGB identifies significantly more true positives than other methods with false discovery control at lower target FDRs.
+For example, in the regression results in Figure 1, IPSSGB identifies 70% of important features when the target FDR is 0.1, while IPSSRF identifies 40% and the remaining methods identify close to none.
+IPSSGB’s TPR even surpasses the TPR of methods without error control in almost all experiments despite having far fewer false positives.
+With an average runtime of less than 15 seconds in all experiments, IPSSGB is also one of the fastest methods.
 Among the other methods with theoretical error control, IPSSRF performs well in terms of identifying true positives while controlling false positives, though not as well as IPSSGB.
-IPSSL, whose model assumptions are violated, performs poorly, neither controlling false positives at the target E(FP) or FDR levels, nor identifying many true ones.
-SSBoost, whose curves are obtained by varying the target E(FP) between 0 and 10, has few false positives but also few true positives.
-Observe, for example, that the for SSBoost in Figure 3 is always significantly lower than those for IPSSGB and IPSSRF, despite all three methods having the same target E(FP).
-This is due in part to the weakness of the efp scores used by SSBoost relative to those used by IPSS (Melikechi and Miller, 2024).
-KOBT performs poorly in general; most notably, its actual FDRs far exceed its target FDRs of 0.1 and 0.3 in all of the simulation settings in Figure 4.
-KOBT also runs in around 90 seconds, making it one of the slowest methods.
-With the exception of IPSSL for classification, all of the stability selection-based methods run in at most 40 seconds, often less.
-The absence of error control for Boruta and Vita is clearly apparent in Figures 3 and 4.
-Both methods usually have FDRs over 0.75—often 20 or more false positives on average—far surpassing IPSSGB and IPSSRF.
-Despite this, Boruta and Vita identify fewer true positives than IPSSGB with a target E(FP) of 3 and, in many cases, IPSSGB with a target FDR of 0.3.
-The RF and XGBoost curves are obtained by increasing the number of features selected, as ranked by their importance scores, from 0 to 30; that is, the first nonzero point along the curve is obtained by keeping only the feature with the highest importance score, the second point is obtained by keeping the features with the two largest importance scores, and so on.
-Interestingly, RF and XGBoost significantly underperform IPSSGB and IPSSRF in terms of true and false positives across the whole 0 to 30 range, indicating that IPSS is able to identify important features that simple thresholding of the baseline algorithm cannot, no matter how many features are kept.
-In terms of runtime, RF and XGBoost run quickly except in the $p=5000$ regression experiment.
-Boruta and Vita are typically slower than the IPSS methods, and this disparity grows with the number of features.
+IPSSL1, whose parametric assumptions are violated, performs poorly, failing to control false positives at target FDR levels.
+It also identifies essentially no true positives in the Gaussian experiments.
+SSBoost is overly conservative, undershooting the target FDR at the expense of identifying few true positives.
+This is partly due to the weakness of the efp scores used by SSBoost relative to those used by IPSS (Melikechi and Miller, 2024).
+All versions of model-X knockoffs underperform IPSSGB and IPSSRF in all experiments.
+This is even true in the Gaussian setting, where the known joint distribution of the features was used to implement these methods (recall that model-X knockoffs requires knowledge of the joint distribution, while IPSS does not).
+With the exception of KOL and KORF in the $p=5000$ and, to a lesser extent, $p=2000$ RNA-seq experiments, all of these methods control the FDR at target levels.
+KORF is the only knockoffs-based method that identifies essentially any true positives in the Gaussian experiments.
+DeepPINK, which combines knockoffs with deep neural networks, has the lowest TPR out of every method in almost all experiments.
+This agrees with our earlier observation that deep learning typically underperforms tree-based methods on tabular data and requires tens of thousands of samples for competitive performance (Grinsztajn et al., 2022).
+The absence of error control for Boruta and Vita is clearly apparent in Figures 2 and S2.
+Both methods usually have FDRs over 0.75, far surpassing other methods.
+Despite this, Boruta and Vita almost always identify fewer true positives than IPSSGB when the target FDR is greater than 0.2 or, in some cases, even 0.1.
+Boruta and Vita are also slower than other methods, and this disparity grows with the number of features (Tables S3 and S4).
+VSURF usually has a lower FDR than Boruta and Vita, but still underperforms IPSSGB.
+Its excessive runtimes prevented us from including it in the $p=2000$ and $5000$ RNA-seq experiments.
+Several other methods are also omitted, namely KOBT, RFEGB, and RFHT.
+Briefly, KOBT far exceeded target FDRs despite extensive tuning, while RFEGB and RFHT had FDRs over $0.8$ and extremely long runtimes.
+For details, see Sections S3 and S2.
 Figure 2.
-TP versus FP for the six simulation experiments.
-Columns correspond to the number of features, $p$, and the top and bottom rows correspond to regression and classification, respectively.
-Curves for the three IPSS methods and SSBoost are obtained by varying the target E(FP), with showing results when the target E(FP) is 3.
-The KOBT curve is obtained by varying the target FDR, and the RF and XGBoost curves are obtained by increasing the number of features selected (as ranked by their importance scores) from 0 to 30.
-The remaining methods, indicated by the symbols, do not have curves because they select fixed sets of features.
+RNA-seq simulation results (regression).
+First, second, and third columns correspond to the $p=500$, $2000$, and $5000$ experiments, respectively.
+The horizontal axis in each plot shows the target FDR.
+Methods without false discovery control (gray horizontal lines) do not vary with the target FDR.
+The black dashed line in the FDR plots represents perfect FDR calibration, FDR = target FDR.
+Non-gray solid lines correspond to IPSS or stability selection-based methods.
+Non-gray dashed lines show methods based on model-X knockoffs.
+4.  Cancer studies
+We study ovarian cancer and glioma (a type of brain cancer).
+Both studies include multiple substudies in which the features are either genes, measured by RNA-seq, or microRNAs (miRNAs).
+The response variables are either prognosis (whether the patient was alive at last follow-up), tumor purity (the proportion of cancerous cells in a tissue sample), or the expression level of a particular gene or miRNA.
+All data are from The Cancer Genome Atlas (Weinstein et al., 2013) and were downloaded from LinkedOmics (Vasaikar et al., 2018).
+Additional details and results are in Section S4.
+We assess feature selection performance by (i) performing literature searches and (ii) assessing predictive performance using cross-validation (see Sections S4.3 and S4.4, respectively).
+The literature search results show that IPSSGB and IPSSRF consistently identify more known important features at lower target FDRs than other methods.
+In Table S6, for example, IPSSGB and IPSSRF identify 8 and 6 miRNAs, respectively, all but one of which have been implicated in ovarian cancer prognosis.
+In contrast, IPSSL1 identifies 4 miRNAs, missing the three most significant ones, while KOGLM, KORF, KOL1, DeepPINK, and SSBoost select no miRNAs at all, even when the target FDR is $0.5$.
+In the RNA-seq and glioma prognosis study (Table S10), only IPSSGB identifies the key oncogene FOXM1 (Raychaudhuri and Park, 2011), and only IPSSGB, IPSSRF, and KORF identify WEE1, which is also known to play a significant role in glioma outcomes (Music et al., 2016).
+Furthermore, IPSSGB and IPSSRF are more confident in their selections, assigning WEE1 $q$-values of 0.10 and 0.06, respectively, while KORF does not select WEE1 until the target FDR is reduced to 0.44.
+More generally, Table S10 shows that IPSSGB, IPSSRF, and, to a lesser extent, IPSSL1, tend to identify genes supported by the glioma literature while avoiding genes with little or no known connection.
+In contrast, KORF and especially KOL1 select many genes with limited or no literature support, while KOGLM, DeepPINK, and SSBoost select no genes at all.
+Briefly, our cross-validation (CV) studies proceed as follows.
+In each of 20 CV steps, one group of patients is set aside (the test set), and a set of features is selected by each method using the data in the remaining groups (the training set).
+Next, for each method we construct three predictive models—a linear model, a random forest model, and a gradient boosting model—using only the features selected by that method on the training data.
+Each model is then used to predict responses from the test set, and the smallest of the three prediction errors is recorded (we use mean squared error for regression and $1-\mathrm{accuracy}$ for classification).
+All three models are implemented so that no method has an inherent advantage over another.
+For example, features selected by IPSSL1 may be better suited to linear model predictions than those selected by IPSSGB, while those selected by IPSSGB may be better suited to gradient boosting predictions than the ones selected by IPSSL1.
+Figure 3 shows the results from our RNA-seq and glioma prognosis CV study.
+On average, the top 20 genes selected by IPSSGB yield the same prediction error as the full model that uses all $10,058$ genes.
+IPSSRF and IPSSL1 achieve the smallest prediction errors overall, and only use 10 to 20 genes to do so.
+KORF and KOL1 have higher predictive errors despite using more than 40 selected genes, and Boruta, which selects over 100 genes, has an average error similar to that of the full model.
+Vita (not shown) selects over 500 genes on average and has an average prediction error of 0.22, and DeepPINK and SSBoost select no genes at all, even when the target FDR is 0.5.
 Figure 3.
-TPR versus FDR for the six simulation experiments.
-Results are similar to those in Figure 3, except now we show the FDR and TPR rather than E(FP) and E(TP), and curves for the three IPSS methods and KOBT are obtained by varying the target FDR, with and showing results when the target FDR is 0.1 and 0.3.
-Figure 5.
-Runtimes for the six experiments.
-Columns correspond to the number of features, $p$, and the top and bottom rows correspond to regression and classification, respectively.
-Orange lines denote median runtime and whiskers extend to 1.5 times the interquartile range.
-5.  Ovarian cancer
-We apply IPSSGB, IPSSRF, and the other methods in Table 1 to microRNA (miRNA) data from the cohort of ovarian cancer patients described in Section 4 (Vasaikar et al., 2018).
-We report two studies here; see Section S3 for results from additional studies, including analyses of the RNA-seq data from this cohort.
-In the first study, the features are the 588 miRNAs and the response variable is prognosis, which is defined as whether the patient was still alive at the time of last follow-up.
-In the second study, the response is one of the miRNAs, miR-150, identified as important in the first stage, and the features are the remaining miRNAs.
-Thus, since miRNA measurements are continuous, the first study is a binary classification problem, and the second study is regression.
-All analyses, both here and in Section S3, are conducted using 10-fold cross-validation (CV) as follows.
-First, all patients with missing values are removed, resulting in $n=442$ and $n=453$ for the first and second studies, respectively.
-The remaining patients are evenly partitioned into 10 groups.
-In each of the 10 CV steps and for each selection method, one group of patients is set aside (the test set), and a set of miRNAs is selected by applying the method to the data in the remaining groups (the training set).
-For the IPSS methods, we also compute the efp scores and $q$-values of each miRNA on the training set.
-Next, for each method we construct three predictive models—a linear model (either lasso or $\ell^{1}$-regularized logistic regression), a random forest model, and a gradient boosting model—using only the miRNAs selected by that method.
-Each model is then used to predict responses from the test set, and the smallest of the three prediction errors is recorded.
-We implement all three models so that no method has an inherent advantage over another.
-For example, the miRNAs selected by IPSSL may be better suited to minimizing error in a linear model than those selected by IPSSGB, while the miRNAs selected by IPSSGB may be better suited to minimizing error in a gradient boosting model than those selected by IPSSL.
-The linear and random forest predictive models are implemented with scikit-learn (Pedregosa et al., 2011) and gradient boosting with XGBoost (Chen and Guestrin, 2016), always with default parameters.
-For continuous responses, in each CV step we subtract the mean of the training responses from all responses, training and test, and scale all responses by the empirical standard deviation of the training responses.
-Table 2 shows the top five miRNAs related to prognosis for each IPSS method, as ranked by their median $q$-values over the 10 CV steps.
-In Section S3, we further investigate the findings in Table 2 via a literature review.
-Notably, we find numerous studies that link several of the top miRNAs selected by IPSSGB and IPSSRF, namely miR-30d, miR-93, miR-96 and miR-150, to ovarian cancer prognosis.
-Among these, IPSSL only ranks miR-96 in its top five, and we found little literature identifying its top-ranked feature, miR-1-2, to ovarian cancer.
-The results of the second study, identifying miRNAs that are related to miR-150, are shown in Figure 6.
-Here, IPSSGB achieves the lowest average prediction error over the 10 cross-validation steps of any method, while using just 15 to 20 miRNAs, achieving lower prediction error than even the method that uses all 587 miRNAs to predict miR-150 levels (the dashed gray line in Figure 6).
-Figure 7 is similar to Figure 6, except now the response is tumor purity, the proportion of cancerous cells in a tissue sample.
-Here, we see IPSSGB performs best with the fewest features, while the top 30 features selected by IPSSRF achieve essentially identical predictive performance to the method that uses all 588 miRNAs.
-IPSSGB IPSSRF IPSSL
-miR-96(0.2) miR-1270(0.27) miR-1-2(0.07)
-miR-150(0.28) miR-93(0.36) miR-1270(0.08)
-miR-1270(0.43) miR-30d(0.36) miR-96(0.09)
-miR-30d(0.43) miR-342(0.36) miR-342(0.09)
-miR-1301(0.46) miR-150(0.41) miR-934(0.12)
-Table 2. MicroRNA and prognosis. The top five miRNAs related to prognosis according to each IPSS method, as ranked by their median $q$-values over the 10 cross-validation steps.
-Figure 6.
-MicroRNA and miR-150.
-The horizontal and vertical axes show the average number of features selected and the average prediction error over the 10 cross-validation steps, respectively.
-Curves for the IPSS methods and SSBoost are obtained by varying the target E(FP) between 1/8 and 10.
-Points for these methods correspond to target E(FP)s of 1/8, 1/4, 1/2, 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, with larger target E(FP)s generally corresponding to more features selected and lower prediction error.
-Similarly, points for KOBT correspond to target FDRs of 0.01, 0.025, 0.05, 0.1, 0.15, 0.2, 0.25, 0.3, and 0.4.
-The points for RF and XGBoost correspond to selecting between 1 and 30 features (hence, the number of features selected by these two methods is deterministic).
-Boruta and Vita do not have error control parameters and are therefore indicated by single points rather than curves.
-Figure 7.
-MicroRNA and tumor purity.
-This is the same as Figure 6, except now the response is tumor purity, the proportion of cancerous cells in a sample.
-We again see the IPSS methods outperform other methods, achieving similar or better predictive performance with relatively few features.
+RNA-seq and glioma prognosis.
+The horizontal and vertical axes show the average number of genes selected and the average prediction error over the 20 cross-validation steps, respectively.
+Curves for each method are obtained by varying the target FDR between 0 and 0.5.
+DeepPINK and SSBoost selected no genes and are therefore shown by single points rather than curves.
+Boruta is also indicated by a single point because it does not have FDR control parameters.
+The dashed black line shows the average error when using all $p=10,058$ genes to predict prognosis.
 5.  Discussion
-We have demonstrated that IPSSRF and IPSSGB achieve superior results in terms of false positive control, detecting true positives, and computation time.
+We have demonstrated that IPSSRF and IPSSGB achieve superior results in terms of false positive control, true positive detection, and computation time.
 More broadly, IPSS for thresholding is a general framework whose theory and implementation apply to arbitrary importance scores.
 For instance, examples of other scores include $p$-values (with smaller $p$-values indicating greater importance), Shapley values (used to quantify the contribution of individual features to neural networks and other machine learning models), and loadings in principal components analysis (which quantify the contribution of each feature to a given principal component).
-The main practical limitation to consider is the cost of computing the relevant importance scores, since IPSS must compute these scores on many random subsamples of the data.
-We have also introduced efp scores and shown that, in addition to controlling E(FP), they can control the FDR and estimate $q$-values.
+The main practical limitation to consider is the cost of computing the relevant importance scores, since IPSS must compute these scores on multiple subsamples of the data.
+We have also introduced efp scores and shown that, in addition to controlling E(FP), they can be used to control the FDR and estimate $q$-values.
 Storey (2003) showed that $q$-values admit a Bayesian interpretation, suggesting a link between IPSS and Bayesian feature selection that could be an interesting line of future work.
-It would also be interesting to study whether $q$-values estimated from efp scores are more accurate than $q$-values estimated from other quantities, such as $p$-values (Storey and Tibshirani, 2003).
 Finally, a more ambitious goal is to extend IPSS to unsupervised feature selection problems (that is, feature selection when there is no response variable) and non-iid data.
 The PCA-based scores mentioned above provide at least one way to apply IPSS in an unsupervised setting.
-Developing a rigorous approach to IPSS for non-iid data could provide novel methods for nonparametric feature selection with error control for time series, networks, and longitudinal data.
-Data availability and code
-All of the code and data used in this work are freely available at https://github.com/omelikechi/ipss.
-A Python package for implementing integrated path stability selection, including IPSSGB and IPSSRF, is freely available at https://pypi.org/project/ipss/
+Developing a rigorous approach to IPSS for non-iid data could provide novel methods for nonparametric feature selection with false discovery control for networks and spatially or temporally-indexed data.
+Code and data
+All code and data used in this work are available on GitHub (https://github.com/omelikechi/ipss_bioinformatics) and permanently archived on Zenodo (https://doi.org/10.5281/zenodo.15335289).
+A Python package for implementing IPSS is available on GitHub (https://github.com/omelikechi/ipss) and PyPI (https://pypi.org/project/ipss/).
+An R implementation of IPSS is also available on GitHub (https://github.com/omelikechi/ipssR).
 Acknowledgements
@@
 was supported in part by the National Institutes of Health (NIH) grant R01CA240299.
-Supplementary material for “Fast nonparametric feature selection with error control using integrated path stability selection”
-S1.  Additional IPSS details
-We provide an algorithm that implements integrated path stability selection (IPSS) for thresholding (Section S1.1), elaborate on the theory of IPSS and its connection to efp scores (Section S1.2), and describe and study the IPSS parameters in greater detail (Section S1.3).
-S1.1.  Algorithm.
-Algorithm S1, discussed in Section 2.3, implements IPSS for thresholding.
+SUPPLEMENTARY MATERIALS
+We provide further details about IPSS (Section S1), describe the feature selection methods considered in this work (Section S2), and present results from our simulation studies (Section S3), cancer studies (Section S4), and sensitivity analyses of the IPSS parameters (Section S5).
+S1.  IPSS details
+We provide an algorithm that implements integrated path stability selection (IPSS) for feature importance scores (Section S1.1), elaborate on the theory of IPSS and its connection to efp scores (Section S1.2), and describe the IPSS parameters in greater detail (Section S1.3).
+S1.1  Algorithm.
+Algorithm S1, discussed in Section 2.2, implements IPSS for feature importance scores.
 The number of grid points used to evaluate the integrals in Algorithm S1 is always $K=100$.
 Like many of the other IPSS parameters, $K$ is inconsequential provided it is sufficiently large; in our experience, values greater than 25 suffice.
-This is because the function $f(x)=(2x-1)^{3}\operatorname{\mathds{1}}(x\geq 0.5)$, the paths $\lambda\mapsto\hat{\pi}_{j}(\lambda)$ (which are monotonically increasing functions of $\lambda$), the quantity $\mathcal{I}(\Lambda)$, and the measure $\mu$ are all very numerically stable.
-Algorithm S1 Integrated path stability selection for thresholding
-1:Data $Z_{1:n}$, importance function $\Phi$, number of grid points $K$, and number of iterations $B$.
+This is because the function $f(x)=(2x-1)^{3}\operatorname{\mathds{1}}(x\geq 0.5)$, the paths $\lambda\mapsto\hat{\pi}_{j}(\lambda)$ (which are monotonically increasing functions of $\lambda$), the quantity $\mathcal{I}(\Lambda)$, and the family of measures $\mu_{\delta}(d\lambda)=z^{-1}_{\delta}\lambda^{-\delta}d\lambda$ are all very numerically stable.
+Algorithm S1 Integrated path stability selection for feature importance scores
+1:Data $Z_{1:n}$, importance function $\Phi$, number of grid points $K$ and iterations $B$, probability measure $\mu$, function $f$ (default $f(x)=(2x-1)^{3}\operatorname{\mathds{1}}(x\geq 0.5)$), and cutoff $C$ (default $C=0.05$).
 2:(Optional) Preselect features, as described in Section S2.1.
@@
 9:Initialize $\lambda_{\mathrm{min}}\leftarrow\lambda_{\mathrm{max}}$ and $k\leftarrow 0$.
-10:while $\mathcal{I}([\lambda_{\mathrm{min}},\lambda_{\mathrm{max}}])<0.05$ do
+10:while $\mathcal{I}([\lambda_{\mathrm{min}},\lambda_{\mathrm{max}}])<C$ do
 11:     $\hat{S}_{\lambda}(Z_{A_{b}})=\{j:\Phi_{Z_{A_{b}}}(j)\geq\lambda\}$ for $b=1,\dots,2B$.
@@
 17:$\mathtt{efp}_{Z_{1:n}}(j)=\mathcal{I}(\Lambda)/{\int}_{\Lambda}f(\hat{\pi}_{j}(\lambda))\mu(d\lambda)$ for $j=1,\dots,p$.
-S1.2.  Theory: From IPSS to efp scores.
+S1.2  IPSS theory and efp scores.
 Given the estimated selection probabilities $\hat{\pi}_{j}$, an interval $\Lambda\subseteq(0,\infty)$, a probability measure $\mu$ on $\Lambda$, a function $f:[0,1]\to\mathbb{R}$, and a threshold $\tau$, Melikechi and Miller (2024) define the set of features selected by IPSS by
@@
 Further theoretical and empirical comparisons between Equation S1.4 and other stability selection bounds are available in Melikechi and Miller (2024).
-S1.3.  Parameters.
-As noted above, the function $f(x)=(2x-1)^{3}\operatorname{\mathds{1}}(x\geq 0.5)$ used when implementing IPSS in the main text is determined by the availability and strength of the theoretical bound in Theorem S1.1.
-Similarly, the choice of $\lfloor n/2\rfloor$ samples used to construct the selection probabilities $\hat{\pi}_{j}(\lambda)$ are required for stability selection theorems (not just Theorem S1.1) to hold, and it is unclear how to adapt their proofs to accommodate other sample sizes.
-Thus, $f$ and $\lfloor n/2\rfloor$ are theoretical necessities rather than free parameters.
+S1.3  Parameters.
+Table S1 shows the default IPSS parameters used for IPSSGB and IPSSRF throughout this work (default gradient boosting and random forest parameters are in Section S2.2.1 and Section S2.2.2).
+As noted above, our choice of function $f(x)=(2x-1)^{3}\operatorname{\mathds{1}}(x\geq 0.5)$ is determined by the availability and strength of the theoretical bound in Theorem S1.1.
+Similarly, the choice of $\lfloor n/2\rfloor$ samples used to construct the selection probabilities $\hat{\pi}_{j}(\lambda)$ is required for stability selection theorems (not just Theorem S1.1) to hold, and it is unclear how to adapt their proofs to accommodate other sample sizes.
+Thus, $f$ and $\lfloor n/2\rfloor$ are theoretically determined rather than free parameters.
 The interval $\Lambda=[\lambda_{\mathrm{min}},\lambda_{\mathrm{max}}]$ is determined by setting $\lambda_{\mathrm{max}}$ large enough that no features are selected (see, for example, 7 in Algorithm S1), and setting $\lambda_{\mathrm{min}}$ such that the integral
@@
 in Equation S1.4 is equal to a fixed cutoff $C$.
-As noted in the main text, we always use $C=0.05$, but results are largely independent of this choice (Figures S1 and S2).
-Figures S3 and S4 show IPSS is also indifferent to $\mu$.
-Here, the parameter $\delta$ determines the measure $\mu_{\delta}(d\lambda)=z_{\delta}^{-1}\lambda^{-\delta}d\lambda$, where the normalizing constant $z_{\delta}$ is easily computed in closed form (Melikechi and Miller, 2024).
-The probability measure that we always use in the main text, $\mu_{1}$, averages over $\Lambda$ on a log scale, while $\mu_{0}$ averages on a linear scale.
-The insignificance of $C$ and $\mu$ is unsurprising: Intuitively, the efp scores for IPSS depend primarily on $f$ and the integrand in $\mathcal{I}(\Lambda)$.
+As noted in the main text, we always use $C=0.05$, but results are largely independent of this choice (Figures S12 and S13).
+Figures S14 and S15 show IPSS is also robust to the parameter $\delta$ that determines the measure $\mu_{\delta}(d\lambda)=z_{\delta}^{-1}\lambda^{-\delta}d\lambda$, where the normalizing constant $z_{\delta}$ is easily computed in closed form (Melikechi and Miller, 2024).
+The probability measure $\mu_{1}$ averages over $\Lambda$ on a log scale, while $\mu_{0}$ averages on a linear scale.
+The insignificance of $C$ and $\delta$ is unsurprising: Intuitively, the efp scores for IPSS depend primarily on $f$ and the integrand in $\mathcal{I}(\Lambda)$.
 The actual value of $\mathcal{I}(\Lambda)$ is much less important since the efp scores depend on the relative quantities $\mathcal{I}(\Lambda)/\int_{\Lambda}f(\hat{\pi}_{j}(\lambda))\mu(d\lambda)$ rather than on $\mathcal{I}(\Lambda)$ itself.
-Since $C$ and $\mu$ only affect the value of the bound, not the integrand, they contribute little to the actual performance of IPSS.
-Figure S1.
-Dependence on cutoff (regression).
-IPSSGB is applied to regression data simulated as in Section 4 for cutoff values $C\in\{0.025,0.05,0.075,0.1,0.15\}$, averaged over 100 trials.
-There is little difference between the average true and false positive values for different choices of $C$ over a wide range of target E(FP) values.
-Figures S2, S3, and S4 show similar results for classification and dependence on $\mu$.
-Figure S2.
-Dependence on cutoff (classification).
-IPSSGB is applied to classification data simulated as in Section 4 for cutoff values $C\in\{0.025,0.05,0.075,0.1,0.15\}$, averaged over 100 trials.
-There is little difference between the average true and false positive values for different choices of $C$ over a wide range of target E(FP) values.
-Figure S3.
-Dependence on $\mu$ (regression).
-IPSSGB is applied to regression data simulated as in Section 4 for cutoff values $\delta\in\{0,0.5,1,1.5,2\}$, averaged over 100 trials, where $\delta$ determines to the measure $\mu_{\delta}(d\lambda)=z_{\delta}^{-1}\lambda^{-\delta}d\lambda$.
-There is little difference between the average true and false positive values for different choices of $\delta$ over a wide range of target E(FP) values.
-Figure S4.
-Dependence on $\mu$ (classification).
-IPSSGB is applied to classification data simulated as in Section 4 for cutoff values $\delta\in\{0,0.5,1,1.5,2\}$, averaged over 100 trials, where $\delta$ determines to the measure $\mu_{\delta}(d\lambda)=z_{\delta}^{-1}\lambda^{-\delta}d\lambda$.
-There is little difference between the average true and false positive values for different choices of $\delta$ over a wide range of target E(FP) values.
-S1.4.  Preselection.
-Many feature selection algorithms, including most in Section 3, employ some form of screening, or preselection, as an initial step in the selection process.
-This can be especially helpful in high dimensions for reducing noise and runtime.
-When $p>200$, we preselect features for IPSSGB, IPSSRF, and the other stability selection-based methods (IPSSL and SSBoost, detailed below) by running their respective selection algorithms on the full dataset three times—that is, computing $\Phi_{Z_{1:n}}$ three times—and keeping only the $p_{\text{pre}}=\max\{200,p/20\}$ features with the largest average scores across all three trials.
-Note that $\Phi_{Z_{1:n}}$ is stochastic, given $Z_{1:n}$, due to randomness in the gradient boosting and random forest algorithms.
-This preselection step does not affect the theoretical control on E(FP) since $\lvert\hat{S}_{\mathrm{IPSS},\mathrm{pre}}\cap S^{c}\rvert=\lvert\hat{S}_{\mathrm{IPSS},\mathrm{pre}}\cap S^{c}_{\mathrm{pre}}\rvert$, where $\hat{S}_{\mathrm{IPSS},\mathrm{pre}}$ are the features selected by IPSS using only the preselected features, and $S^{c}_{\mathrm{pre}}$ are the preselected features in $S^{c}$.
-Of course, preselection risks throwing away important features, potentially increasing the number of false negatives.
-However, we have yet to observe a case in which an important feature that was not in the top $p_{\mathrm{pre}}$ preselected features was selected by IPSS when the full set of features was used instead.
-In other words, a feature that does not make the top $p_{\mathrm{pre}}$ is highly unlikely to be selected by IPSS regardless of whether or not preselection is used.
-S2.  Simulation details and additional results
-Algorithm S2 describes the procedure we use to generate data in the simulation study in Section 4.
-See also Figure 1 in the main text for a diagrammatic representation of this process.
+Since $C$ and $\mu$ only affect the value of the bound, not the integrand, they contribute little to the actual performance of IPSS, as indicated by the sensitivity analyses in Section S5.
+Method $B$ $C$ $f(x)$ $\delta_{\text{reg}}$ $\delta_{\text{class}}$
+IPSSGB $100$ $0.05$ $(2x-1)^{3}\operatorname{\mathds{1}}(x\geq 0.5)$ $1.25$ $1$
+IPSSRF $50$ $0.05$ $(2x-1)^{3}\operatorname{\mathds{1}}(x\geq 0.5)$ $1.25$ $1.25$
+Table S1.
+Default IPSS parameters.
+Both IPSSGB and IPSSRF always use $C=0.05$ and $f(x)=(2x-1)^{3}\operatorname{\mathds{1}}(x\geq 0.5)$.
+IPSSRF uses $B=50$ to reduce runtimes without any noticeable difference in selection performance.
+The parameters $\delta_{\text{reg}}$ and $\delta_{\text{class}}$ determine the measure $\mu_{\delta}$ in regression and classification problems, respectively.
+S2.  Overview of methods and implementation details
+We describe the preselection procedure used to improve feature selection (Section S2.1) and provide implementation details for each feature selection method considered in this work (Section S2.2).
+S2.1  Preselection.
+Many feature selection methods employ some form of screening, or preselection, as an initial step in the selection process.
+This can be especially helpful in high dimensions for increasing power and reducing runtimes.
+For many of the methods in Section S2.2, we preselect features by running a randomized importance function on the full dataset three times—that is, computing $\Phi_{Z_{1:n}}$ three times—and keeping only the $p_{\text{pre}}$ features with the largest average scores across all three trials.
+For example, preselection for IPSSRF entails fitting three random forests to the full dataset and keeping the $p_{\text{pre}}$ features with the largest average importance scores.
+For IPSS, this preselection step does not affect the theoretical control on E(FP) since $\lvert\hat{S}_{\mathrm{IPSS},\mathrm{pre}}\cap S^{c}\rvert=\lvert\hat{S}_{\mathrm{IPSS},\mathrm{pre}}\cap S_{\mathrm{pre}}^{c}\rvert$, where $\hat{S}_{\mathrm{IPSS},\mathrm{pre}}$ are the features selected by IPSS using only the preselected features, and $S_{\mathrm{pre}}^{c}$ are the preselected features in $S^{c}$.
+Of course, preselection risks discarding important features, potentially increasing the number of false negatives.
+In practice, however, we find that preselection actually helps IPSS, stability selection, and model-X knockoffs identify more true positives while still controlling false discoveries.
+This is because preselection gets rid of many noisy features, making it easier for these methods to detect the true signal.
+In Section S2.2, we describe the preselection parameters for each method, which were determined by extensive testing on simulated data.
+We do not apply preselection when implementing Boruta, RFEGB, RFHT, Vita, and VSURF since these methods include their own internal screening steps.
+S2.2  Methods.
+We describe how each method in Table S2 is implemented in this work.
+Method Package Error control Base method Non-default settings
+IPSSGB ipss (Python) Boosting
+IPSSRF ipss (Python) Random forest
+IPSSL1 ipss (Python) Lasso
+KOGLM knockoff (R) GLM
+KOL1 knockoff (R) Lasso
+KORF knockoff (R) Random forest
+DeepPINK knockpy (Python) Neural network
+SSBoost XGBoost (Python) Boosting $\texttt{assumption}=r\mathrm{-concave}$
+with stabs (R) $\tau=0.75$
+KOBT KOBT (R) Boosting $\texttt{num}=100$, $\texttt{bound}=200$,
+[…]
+RFHT rfvimptest (R) Random forest
+Boruta Boruta (R) Random forest
+RFEGB scikit-learn (Python) Boosting
+Vita vita (R) Random forest $p$-value threshold $=0$
+VSURF VSURF (R) Random forest VSURF_pred
+Table S2.
+Feature selection methods.
+Software packages are listed with the language used to implement them in parentheses.
+Details about each method, including descriptions of their non-default settings, are in Sections S2.2.1S2.2.7.
+For methods with no non-default settings, we use the default settings in their respective packages.
+S2.2.1  IPSSGB.
+The IPSS-related parameters used to implement IPSSGB are in Table S1.
+For preselection, we use gradient boosting as the baseline selection algorithm and set $p_{\text{pre}}=100$.
+We implement gradient boosting using XGBoost (Chen and Guestrin, 2016).
+All XGBoost parameters are set to their default values except for two changes: The proportion of features considered when splitting each node (called colsample_bynode in XGBoost and often $\mathtt{mtry}$ elsewhere) is changed from $1$ to $1/3$, and the maximum depth of each tree (max_depth) is changed from $6$ to $1$, making each tree a stump.
+The latter change significantly improved the performance of IPSSGB, both in terms of speed and feature selection.
+S2.2.2  IPSSRF.
+The IPSS-related parameters used to implement IPSSRF are in Table S1.
+For preselection, we use random forests as the baseline selection algorithm and set $p_{\text{pre}}=100$.
+We implement random forests using scikit-learn (Pedregosa et al., 2011).
+All random forest parameters are set to their default values except for two changes: The proportion of features considered when splitting each node (called max_features in scikit-learn) is changed from $1$ to $1/10$, and the number of trees used to build each random forest (n_estimators) is changed from $100$ to $50$.
+These changes improved the efficiency of IPSSRF without sacrificing its feature selection performance.
+S2.2.3  IPSSL1.
+As discussed in the main text, IPSSL1 is a parametric version of IPSS based on $\ell^{1}$-regularization (Melikechi and Miller, 2024).
+For regression, the baseline algorithm is lasso (Tibshirani, 1996), and for classification, it is $\ell^{1}$-regularized logistic regression (Friedman et al., 2010).
+All parameters are set to their default values in the ipss Python package: https://pypi.org/project/ipss/.
+S2.2.4  Model-X knockoffs (KOGLM, KOL1, KORF, DeepPINK, KOBT).
+Model-X knockoffs work by first constructing knockoffs $\tilde{X}=(\tilde{X}_{1},\dots,\tilde{X}_{p})$ of the original features $X=(X_{1},\dots,X_{p})$.
+By definition, $\tilde{X}$ must satisfy (i) the joint distribution of $(X,\tilde{X})$ is invariant under pairwise exchanges of the original features $X_{j}$ and their corresponding knockoffs $\tilde{X}_{j}$, and (ii) $\tilde{X}$ is conditionally independent of the $Y$ given $X$ (see Definition 2 in Candes et al.
+(2018)).
+Once knockoffs are constructed, feature importance scores—called feature statistics in the knockoffs literature—are computed for all of the original and knockoff features and subsequently used to select original features in a way that controls the FDR.
+Like IPSS, any feature importance function can be used.
+As noted in the main text, model-X knockoffs requires knowledge of the joint distribution of $X$.
+In our multivariate Gaussian simulations in Section 3, all of the model-X knockoffs methods are implemented using the true, known joint distribution.
+In the remaining examples, where the joint distribution of $X$ is not known, we use the default methods for constructing approximate model-X knockoffs (for example, second-order Gaussian knockoffs in the knockoff R package).
+KOGLM, KOL1, and KORF are all implemented using the R package knockoff.
+KOGLM uses feature importance scores from a generalized linear model (GLM); we find it performs best when using random forests for preselection with $p_{\text{pre}}=200$.
+KOL1 uses feature importance scores from lasso (regression) or $\ell^{1}$-regularized logistic regression (classification).
+Like IPSSL1, we use lasso (or logistic regression) for preselection, with $p_{\text{pre}}=200$.
+KORF uses feature importance scores from random forests; like KOGLM, we find it performs best when using random forests for preselection with $p_{\text{pre}}=200$.
+DeepPINK uses deep neural networks to construct feature importance scores. We implement it using the Python package knockopy and random forests for preselection with $p_{\text{pre}}=100$.
+KOBT is implemented with the R package KOBT.
+It uses boosted tree models to construct importance scores.
+We tested KOBT numerous times on simulated data with many different tuning and preselection parameters (including no preselection), but found that KOBT consistently and dramatically exceeded its target FDR.
+For this reason, we omit its results from all plots and tables in this work.
+Finally, we tested all of the model-X knockoffs methods without using preselection. In these cases, power was often extremely low and the runtimes were much longer.
+S2.2.5  SSBoost.
+The closest method to IPSSGB in terms of its underlying approach is that of Hofner et al.
+(2015), referred to here as SSBoost.
+Unlike IPSSGB—which uses importance scores from gradient boosting—SSBoost applies stability selection to choose the number of features used per boosting run.
+Furthermore, IPSSGB uses IPSS to construct efp scores, whereas SSBoost uses a version of stability selection introduced by Shah and Samworth (2013).
+This is perhaps the most significant difference since the efp scores for IPSS have much tighter bounds than those for other forms of stability selection (Melikechi and Miller, 2024).
+From a practical standpoint, this causes other versions of stability selection to identify fewer important features than IPSS.
+Hofner et al.
+(2015) provide code for SSBoost that combines the R packages mboost (Hofner et al., 2014) and stabs (Hofner and Hothorn, 2017) in the form of a worked example, but the mboost implementation of boosting was prohibitively slow in the dimensions we consider.
+By adapting their code to use XGBoost in place of mboost for boosting and by preselecting features using gradient boosting with $p_{\text{pre}}=150$, we were able to reduce SSBoost runtimes considerably with no apparent change in results.
+The XGBoost parameters used for SSBoost are the same as those used for IPSSGB (Section S2.2.1).
+For the stability selection part of SSBoost, we use the default parameters in stabs, and the selection threshold is set to $\tau=0.75$, which is the middle of the interval $(0.6,0.9)$ recommended by Meinshausen and Bühlmann (2010).
+S2.2.6  RFHT.
+We tested RFHT (Coleman et al., 2022), which achieves theoretical error control by using random forests for hypothesis testing.
+However, one test run with default parameters on simulated data with 500 samples, 500 features, and 20 true features took over 51 minutes, returning 15 true positives and 43 false positives.
+By contrast, IPSSGB with a target FDR of $0.2$ took 11 seconds and returned 10 true positives and 0 false positives on the same data.
+This method is omitted from our studies because its performance does not appear to justify its excessive runtime.
+S2.2.7  Methods without false discovery control (Boruta, RFEGB, Vita, VSURF).
+Boruta, Vita, and VSURF are implemented using R packages of the same names.
+Boruta is run with default parameters.
+For Vita, we set the $p$-value threshold to 00.
+For VSURF, we use the function VSURF_pred rather than VSURF_interp to select the final set of features (Genuer et al., 2010).
+Both of these choices favor sparsity, which aligns well with our simulation designs.
+As noted in the main text, VSURF is too computationally expensive to include in our $p=2000$ and $5000$ simulation studies; see also Table S3.
+We implement RFEGB by combining XGBoost and scikit-learn.
+On simulated Gaussian data with $n=250$ and $p=500$, one run of RFEGB took over 12 minutes when removing five features per iteration (the default is one feature removed per iteration, which takes approximately 5 times as long).
+For comparison, IPSSGB ran in 5 seconds on the same data and had far fewer false positives and similar power.
+Due to its high computational cost, poor performance, and many tuning parameters (which is true of recursive feature elimination in general), RFEGB is largely omitted from this work.
+S3.  Simulation results and details
+We present additional simulation results (Section S3.1) and RNA-seq simulation details (Section S3.2).
+S3.1  Additional simulation results.
+Figure S1 shows the $n=250$ multivariate Gaussian simulation results, described in Section 3.2.
+Figure S2 shows the $p=500$, $2000$, and $5000$ RNA-seq simulation results for classification, described in Section 3.3.
+Tables S3 and S4 show the average runtimes of each method in each simulation experiment.
+Figure S1. Gaussian simulation results ($n=250$). See Figure 1 for details.
+Figure S2. RNA-seq simulation results (classification). See Figure 2 for details.
+Method MG($n=250$) MG($n=500$) RNA($p=500$) RNA($p=2000$) RNA($p=5000$)
+IPSSGB 4.4 (0.11) 6.3 (0.01) 6.7 (0.23) 9.0 (0.43) 14.6 (0.45)
+IPSSRF 4.5 (0.05) 7.5 (0.10) 7.2 (0.14) 10.8 (0.29) 19.6 (0.63)
+IPSSL1 1.6 (0.23) 1.0 (0.05) 4.0 (0.91) 3.1 (0.66) 6.4 (0.76)
+KOGLM 10.8 (0.92) 12.4 (0.91) 10.0 (1.92) 12.6 (1.50) 19.2 (1.66)
+KOL1 11.2 (0.88) 12.0 (1.04) 7.5 (1.27) 6.7 (0.93) 7.2 (1.16)
+KORF 12.8 (1.16) 14.3 (1.01) 4.2 (0.50) 7.4 (0.22) 13.9 (0.50)
+DeepPINK 5.0 (3.55) 6.0 (2.79) 5.9 (4.17) 8.8 (0.21) 15.4 (0.42)
+SSBoost 4.1 (0.04) 6.7 (0.03) 6.7 (0.04) 8.1 (0.15) 13.7 (0.08)
+Boruta 42.2 (3.09) 120.7 (7.93) 42.0 (5.96) 57.0 (6.51) 84.5 (6.66)
+Vita 9.5 (0.19) 24.2 (0.40) 20.6 (0.47) 83.5 (4.09) 195.3 (4.34)
+VSURF 196.8 (8.16) 548.6 (23.68) 286.7 (72.36)
+RFHT 3087*
+RFEGB 749*
+Table S3.
+Average runtimes (in seconds) over 100 trials for each regression experiment.
+MG stands for Multivariate Gaussian.
+Standard deviations are in parentheses.
+Recall that $p=500$ in the MG experiments and $n=500$ in the RNA experiments, and that VSURF was too computationally expensive to include when $p=2000$ and $5000$As discussed in Sections S2.2.6 and S2.2.7, RFHT and RFEGB are largely omitted due to their poor selection performance and excessive runtimes in initial tests (shown in the table).
+Hence, their remaining entries are blank.
+Method MG($n=250$) MG($n=500$) RNA($p=500$) RNA($p=2000$) RNA($p=5000$)
+IPSSGB 4.3 (0.04) 6.5 (0.11) 6.6 (0.07) 7.5 (0.20) 14.3 (0.42)
+IPSSRF 4.2 (0.02) 7.0 (0.05) 6.8 (0.10) 9.9 (0.22) 17.7 (0.49)
+IPSSL1 5.9 (0.15) 6.0 (0.22) 8.4 (0.86) 9.3 (1.17) 13.1 (1.25)
+KOGLM 10.6 (0.10) 12.0 (0.14) 9.3 (1.57) 11.8 (1.67) 17.4 (1.64)
+KOL1 10.7 (0.12) 11.9 (0.21) 7.2 (1.13) 6.0 (0.83) 6.3 (0.80)
+KORF 10.1 (0.95) 12.1 (0.17) 3.1 (0.10) 5.9 (0.18) 11.3 (0.38)
+DeepPINK 4.7 (0.29) 5.7 (0.28) 5.2 (0.10) 8.0 (0.24) 13.3 (0.43)
+SSBoost 4.0 (0.01) 6.6 (0.04) 6.7 (0.02) 8.2 (0.04) 13.8 (0.11)
+Boruta 27.8 (1.81) 76.6 (6.29) 25.4 (3.91) 37.4 (4.51) 58.6 (3.99)
+Vita 4.8 (0.06) 10.7 (0.06) 9.2 (0.16) 35.3 (0.72) 89.1 (4.59)
+VSURF 78.1 (4.14) 190.9 (8.93) 77.0 (28.08)
+Table S4.
+Average runtimes (in seconds) over 100 trials for each classification experiment.
+MG stands for Multivariate Gaussian.
+Standard deviations are in parentheses.
+Recall that $p=500$ in the MG experiments and $n=500$ in the RNA experiments.
+VSURF was too computationally expensive to include when $p=2000$ and $5000$.
+S3.2  RNA-seq simulation details.
+Algorithm S2 describes the data generating procedure for the RNA-seq simulation studies (Section 3.3), which is also depicted in Figure S3.
 In all steps, “randomly select” means select a parameter uniformly at random from its domain, which are shown in Table S5.
@@
 For example, if the function in the left-most panel of Figure S4 is applied to the genes in a given partition of $S$, then those genes will only significantly affect $Y$ if their collective expression level is positive, while collective expression levels less than $-1$ will have virtually no effect on $Y$.
-Parameter $\alpha$ $\beta$ $\gamma$ $\delta_{1}$ $\delta_{2}$ $p_{\mathrm{true}}$ SNR $u$
-Range $(0.5,1.5)$ $(-1,1)$ $(1,3)$$\{-1,1\}$$\{-1,1\}$$\{10,\dots,30\}$ $(0.5,2)$ $(1,3)$
+Parameter $\alpha$ $\beta$ $\gamma$ $\delta_{1}$ $\delta_{2}$
+Range $(0.5,1.5)$ $(-1,1)$ $(1,3)$$\{-1,1\}$$\{-1,1\}$
 Table S5. Simulation parameters. Parameters are drawn uniformly at random from their corresponding ranges prior to each simulation trial.
 Algorithm S2 Data generation for simulation study (one trial)
-1:RNA-seq data $X_{\mathrm{full}}\in\mathbb{R}^{596\times 6426}$, number of samples $n$, number of features $p$, number of true features $p_{\mathrm{true}}$, signal-to-noise ratio SNR, function parameter domain $\Theta$.
+1:RNA-seq data $X_{\mathrm{full}}\in\mathbb{R}^{596\times 6426}$, number of samples $n$, number of features $p$, number of true features $\lvert S\rvert$, signal-to-noise ratio SNR, function parameter domain $\Theta$.
 2:Randomly select $n$ rows and $p$ columns of $X_{\mathrm{full}}$. Denote the resulting matrix by $X\in\mathbb{R}^{n\times p}$.
 3:Standardize the columns of $X$ to have mean 0 and variance 1.
-4:Randomly select true features $S\subseteq\{1,\dots,p\}$ with $\lvert S\rvert=p_{\mathrm{true}}$.
-5:Randomly select $G\in\{\lfloor p_{\mathrm{true}}/2\rfloor,\dots,p_{\mathrm{true}}\}$. Partition $S$ into $G$ disjoint groups, $S=\bigsqcup_{g=1}^{G}S_{g}$.
+4:Randomly select $\lvert S\rvert$ true features $S\subseteq\{1,\dots,p\}$.
+5:Randomly select $G\in\{\lfloor\lvert S\rvert/2\rfloor,\dots,\lvert S\rvert\}$. Partition $S$ into $G$ disjoint groups, $S=\bigsqcup_{g=1}^{G}S_{g}$.
 6:Initialize the signal, $\eta\leftarrow(0,\ldots,0)^{\mathtt{T}}\in\mathbb{R}^{n}$.
@@
 15:Features $X\in\mathbb{R}^{n\times p}$, responses $y\in\mathbb{R}^{n}$, and important features $S\subseteq\{1,\ldots,p\}$.
-Figure S5. Regression simulation. Empirical E(FP) and E(TP) versus target E(FP) for each IPSS method. The dotted black line shows where E(FP) = target E(FP).
-Figure S6. Classification simulation. Empirical E(FP) and E(TP) versus target E(FP) for each IPSS method. The dotted black line shows E(FP) = target E(FP).
+[…]
+Figure S3.
+Simulation diagram.
+Rows and columns are randomly selected from the full RNA-seq dataset to create a matrix $X$ whose columns are split into important features, $X_{S}$, and unimportant features, $X_{S^{c}}$.
+The columns of $X_{S}$ are further partitioned into $G$ groups; the above figure shows $G=3$.
+The features in each group are summed to obtain $\xi_{g}$, and a different realization $f_{\theta_{g}}$ of $f_{\theta}$ is applied to $\xi_{g}$ for each $g$.
+For regression, response is the sum of the group-specific signals, $f_{\theta_{g}}(\xi_{g})$, plus noise.
 Figure S4. Some realizations of the randomized function $f_{\theta}$.
-One benefit of IPSS is that changing the target E(FP) does not require rerunning the algorithm.
-Figures S5 and S6 show E(TP) and E(FP) as functions of the target E(FP).
-IPSSGB has the best performance in terms of true positives across essentially all target E(FP) values.
-IPSSGB and IPSSRF perform well in the regression setting, and both slightly overshoot the target E(FP) in classification for values less than 6.
-IPSSL performs poorly in general, having relatively few true positives on average and false positives that exceed the target E(FP).
-S3.  Additional ovarian cancer results
-In Section S3.1, we expand upon the ovarian cancer results presented in Section 5 of the main text, namely identifying miRNAs related to prognosis, and miRNAs related to miR-150, respectively.
-In Section S3.2, we perform a similar analysis but for RNA-seq data from the same cohort, identifying genes related to ovarian cancer prognosis, and genes related to one of the genes, AKT2, that was found to be related to prognosis.
-S3.1.  MicroRNA and ovarian cancer.
-We briefly summarize results from the literature that link the miRNAs in Table 2 to ovarian cancer, with a specific focus on prognosis.
-For many of these miRNAs, there are numerous studies connecting them to ovarian cancer beyond what we present.
-Thus, to roughly quantify the relevance of the miRNAs in Table 2 to ovarian cancer, we searched the specific miRNA together with “ovarian cancer” in the Web of Science database, and report both the number of publications and the total number of citations across all returned results.
+Figure S5.
+Distributions of five randomly selected genes from the RNA-seq dataset.
+Features in the ovarian cancer RNA-seq dataset follow a variety of empirical marginal distributions.
+For example, the standardized empirical distributions of the five randomly selected genes above, from left to right, are relatively flat, skewed left, approximately Gaussian, skewed right, and contain outliers.
+Furthermore, the genes exhibit complex correlation structures, with maximum and average absolute pairwise correlations of approximately 0.95 and 0.17 after standardization, respectively.
+S4.  Cancer study results
+We describe our ovarian cancer and glioma studies (Sections S4.1 and S4.2) and present our literature search and cross-validation results (Sections S4.3 and S4.4).
+S4.1  Ovarian cancer.
+For the ovarian cancer cohort, we study the following feature and response combinations (the number of samples and features are shown in parentheses): (i) miRNAs and prognosis ($n=442$, $p=585$), (ii) miRNAs and tumor purity ($n=451$, $p=585$), (iii) miRNAs and miR-150 ($n=453$, $p=584$), (iv) genes and prognosis ($n=549$, $p=6426$), and (v) genes and AKT2 ($n=569$, $p=6425$).
+In all cases, missing values were removed.
+Gene expression levels are measured by RNA-seq.
+We chose miR-150 and the gene AKT2 as responses in studies (iii) and (v), respectively, because literature searches indicated that both are highly related to ovarian cancer.
+Literature search results are only reported for clinical responses (prognosis and tumor purity).
+S4.2  Glioma.
+For the glioma cohort, we study the following feature and response combinations (the number of samples and features in each dataset are shown in parentheses): (i) miRNAs and prognosis ($n=477$, $p=787$), (ii) miRNAs and miR-155 ($n=512$, $p=786$), (iii) genes and prognosis ($n=625$, $p=10,058$), and (iv) genes and FOXM1 ($n=669$, $p=10,057$).
+In all cases, missing values were removed.
+Gene expression levels are measured by RNA-seq.
+The original glioma RNA-seq dataset contains over $20,000$ genes; we only study the roughly $10,000$ genes in the top 50th percentile of average gene expression (discarding lowly expressed genes is common practice).
+We chose miR-155 and the gene FOXM1 as responses in studies (ii) and (iv), respectively, because literature searches indicated that both are highly related to glioma.
+Literature search results are only reported when the response is prognosis.
+S4.3  Literature search.
+In each of the forthcoming studies, we perform literature searches to validate our findings.
+For the microRNA (miRNA) and ovarian cancer prognosis study, we briefly summarize literature supporting each miRNA selected by at least one feature selection method.
+Providing such summaries for every study is beyond the scope of this work.
+Thus, to roughly quantify the relevance of selected features in subsequent studies, we searched the feature name, cancer type, and the word “prognosis” in Europe PMC, an open-access database containing millions of life sciences publications (Europe PMC Consortium, 2015).
+We then report the total number of citations among all articles returned by the search.
+For example, in Table S6, searching “miR-93” + “ovarian cancer” + “prognosis” in Europe PMC returned 43,996 citations.
+For each method, we also include the number of features it selected that had above a certain number of citations, below a certain number of citations, and the total number of features it selected below a certain target FDR.
 We emphasize that these metrics are less important than the papers themselves.
-For example, citation counts favor older publications, and our search criterion does not guarantee that the returned results are relevant to the problem at hand (though the summaries below show that at least some of them are).
-Table S2 is similar to Table 2, except that it shows the total number of publications and citations from our Web of Science literature search and includes some $q$-values that were omitted from Table 2, since Table 2 only reported the five miRNA with the lowest $q$-values for each method.
-Dashed lines indicate that the feature was not identified by the corresponding method at all.
-miRNA Citations Publications IPSSGB IPSSRF IPSSL
-miR-150 1635 26 0.28 0.41 0.2
-miR-93 1270 28 0.36
-miR-30d 596 12 0.43 0.36 0.16
-miR-96 338 10 0.2 0.53 0.09
-miR-342 201 9 0.36 0.09
-miR-1270 175 5 0.43 0.27 0.08
-miR-934 27 2 0.12
-miR-1301 11 3 0.46
-miR-1-2 5 1 0.07
-Table S2. MicroRNA and prognosis. MiRNA are ordered by citations. A missing $q$-value means that miRNA was not selected by the corresponding version of IPSS.
+Citation counts favor older publications, and our search criterion does not guarantee relevance to the specific problem at hand (though the summaries below suggest that at least some results are meaningful).
+Below, we briefly summarize literature relating miRNAs in Table S6 to ovarian cancer. Additional details are available in the associated references.
 miR-1-2. Kandettu et al. (2022) found that miR-1-2 is differentially expressed between cancerous and non-cancerous ovarian cancer cells, but we found no literature linking miR-1-2 to prognosis.
@@
 They conclude that targeting miR-96 is a potentially promising strategy for treating ovarian cancer.
-They also report that miR-96 inhibits phosphorylation of AKT, a gene identified by IPSSGB as being relevant to ovarian cancer prognosis (Section S3.2).
+They also report that miR-96 inhibits phosphorylation of AKT, a gene identified by IPSSGB as being relevant to ovarian cancer prognosis.
 Yang et al.
@@
 The study also showed that miR-342 decreases the expression of key proteins involved in the Wnt/$\beta$-catenin signaling pathway, which may explain its effects on reducing ovarian cancer cell viability and growth.
-miR-934.
-Hu et al.
-(2019) found that miR-934 promotes ovarian cancer cell proliferation and inhibits the death of ovarian cancer cells by targeting the gene BRMS1L, making it is a potential biomarker or therapeutic target for ovarian cancer.
 miR-1270. Ghafouri-Fard et al. (2022) found that miR-1270 plays a role in sensitivity to the chemotherapy drug cisplatin.
@@
 Yu and Gao (2020) found that targeting miR-1301 can inhibit the proliferation of cells that are resistant to the chemotherapy drug cisplatin, thus reducing the occurrence and development of drug-resistant ovarian cancer.
-Tables S3 and S4 show the top ten features for each IPSS method, ranked by their median $q$-values over the 10 CV steps, when the features are miRNA and the responses are miR-150 (Table S3) and tumor purity (Table S4), respectively.
-Tumor purity is the proportion of cancerous cells in a sample.
-Observe that the $q$-values in Table S3 are much smaller than those in Table S2.
-This is to be expected since the signal between miRNAs, in particular, between miR-150 and certain other miRNAs, is likely much stronger than the signal between miRNAs and prognosis.
-IPSSGB IPSSRF IPSSL
-miR-142(0.031) miR-146a(0.016) miR-146a(0.028)
-miR-146a(0.031) miR-142(0.016) miR-142(0.028)
-miR-1247(0.031) miR-223(0.016) miR-1247(0.028)
-miR-642(0.032) miR-642(0.016) miR-139(0.028)
-miR-342(0.036) miR-155(0.016) miR-342(0.029)
-miR-139(0.042) miR-1247(0.016) miR-642(0.029)
-miR-1249(0.051) miR-766(0.016) miR-96(0.029)
-miR-96(0.064) miR-342(0.017) miR-511-1(0.03)
-miR-204(0.072) miR-1228(0.018) miR-34b(0.032)
-miR-34b(0.074) miR-21(0.02) miR-184(0.033)
-Table S3. MicroRNA and miR-150. Top ranked miRNAs for each method in terms of their $q$-values, shown in parentheses.
-IPSSGB IPSSRF IPSSL
-miR-150(0.05) miR-150(0.02) miR-150(0.05)
-miR-22(0.05) miR-22(0.02) miR-22(0.05)
-miR-140(0.08) miR-140(0.02) miR-142(0.05)
-miR-142(0.08) miR-142(0.02) miR-140(0.05)
-miR-1224(0.12) miR-146a(0.03) miR-25(0.06)
-miR-103-2(0.16) miR-223(0.03) miR-1224(0.09)
-miR-223(0.23) miR-214(0.04) miR-135a-2(0.1)
-miR-181c(0.36) miR-132(0.04) miR-155(0.11)
-miR-155(0.49) miR-642(0.05) miR-1247(0.12)
-miR-25(0.52) miR-199b(0.05) miR-34c(0.13)
-Table S4. MicroRNA and tumor purity. Top ranked miRNAs for each method in terms of their $q$-values, shown in parentheses.
-S3.2.  RNA-seq and ovarian cancer.
-In this section, we perform a similar analysis to the miRNA study in Sections 5 and S3.1, but with the 6426 genes from the RNA-seq data described in Section 4.
-Gene Citations Publications IPSSGB IPSSRF IPSSL
-AKT2 9088 118 0.25
-CD38 852 30 0.11 0.31
-PPL 151 8 0.8 0.26
-HOXB4 127 7 0.36
-ZFHX4 105 7 0.78 0.17
-SLAMF7 101 2 0.38 0.36
-WNK1 96 3 0.42 0.23
-SLAMF1 58 5 0.38
-PARD6B 61 1 0.28 0.33
-UGT8 37 1 0.17 0.41
-AP2A1 37 2 0.4
-PROSC 0 0 0.33 0.46 0.2
-MRAP2 0 0 0.09 0.39 0.78
-OR51A4 0 0 0.59 0.2
-CCDC84 0 0 0.2
-SH3BGR 0 0 0.07
+miRNA Citations IPSSGB IPSSRF IPSSL1 KOGLM KORF KOL1 DeepPINK SSBoost
+miR-93 43996 0.35 0.11
+miR-148a 42177 0.39
+miR-150 41195 0.23 0.39
+miR-96 23010 0.23 0.47
+miR-342 20291 0.39 0.23
+miR-30d 19267 0.23 0.33
+miR-301b 3224 0.35
+miR-1270 2543 0.28 0.11 0.21
+miR-1301 1390 0.35
+miR-1-2 1220 0.35 0.21
+$\geq$ 1000 8 6 4 0 0 0 0 0
+$<$ 1000 0 0 0 0 0 0 0 0
+Total 8 6 4 0 0 0 0 0
 Table S6.
-RNA-seq and prognosis.
-Genes are ordered by citations.
-A missing $q$-value means that gene was not selected by the corresponding version of IPSS.
-*A literature search revealed that the gene MRAP2 (also called C6orf117), while not directly related to ovarian cancer, is very much related to obesity, which is linked to poor prognosis and reduced survival times in cancer patients (Chan et al., 2009).
-Figure S8.
-RNA-seq and prognosis.
-No method does well at predicting prognosis relative to the full dataset.
-This is unsurprising, given the many factors that can contribute to poor outcomes which are unlikely to be captured by RNA-seq data alone.
-Nevertheless, several of the genes identified by the various feature selection methods do appear to be related to ovarian cancer prognosis; see Table S5.
-Figure S9. RNA-seq and AKT2. As in several of the other studies, IPSSGB achieves superior predictive performance with fewer features compared to other methods.
-IPSSGB IPSSRF IPSSL
-SUPT5H(0.01) SUPT5H(0.01) SUPT5H(0.02)
-EGLN2(0.01) PAK4(0.01) EGLN2(0.02)
-PAK4(0.01) ADCK4(0.01) ADCK4(0.02)
-ADCK4(0.01) EGLN2(0.01) BLVRB(0.02)
-BLVRB(0.01) ZNF574(0.01) SPSB3(0.03)
-ITPKC(0.02) GSK3A(0.01) HKR1(0.03)
-HKR1(0.04) SARS2(0.01) CAV2(0.04)
-ZMYND17(0.14) ITPKC(0.01) MAP3K10(0.07)
-U2AF2(0.19) CLPTM1(0.01) ACRC(0.08)
-SPSB3(0.2) ZNF526(0.01) ITPKC(0.09)
-Table S6. RNA-seq and AKT2. Top ranked genes for each method in terms of their $q$-values, shown in parentheses.
+MicroRNAs and prognosis (ovarian cancer).
+MiRNAs are ordered by citation count.
+A missing $q$-value indicates the miRNA was assigned a $q$-value of less than 0.5 by the corresponding method.
+The bottom rows report, for each method, the number of selected features with over 1000 citations, under 1000 citations, and the total number selected at the maximum target FDR of 0.5.
+miRNA Citations IPSSGB IPSSRF IPSSL1 KOGLM KORF KOL1 DeepPINK SSBoost
+miR-21 190920 0.08
+miR-155 128839 0.32 0.08 0.09 0.20
+miR-145 89265 0.24
+miR-146a 57373 0.03
+miR-214 57253 0.05
+miR-223 53770 0.24 0.03 0.09
+miR-25 43636 0.12 0.32 0.09 0.20
+miR-22 42211 0.04 0.03 0.07 0.20 0.17
+miR-150 41195 0.04 0.03 0.07 0.20 0.17
+miR-142 39980 0.04 0.03 0.25
+miR-335 33200 0.26
+miR-15b 30988 0.24 0.20 0.25
+miR-140 29667 0.04 0.03 0.17
+miR-152 25767 0.16
+$\geq$ 1000 14 25 12 0 0 7 0 5
+$<$ 1000 0 3 1 0 0 1 0 0
+Total 14 28 13 0 0 8 0 5
+Table S7.
+MicroRNAs and tumor purity (ovarian cancer).
+MiRNAs are ordered by citation count.
+A missing $q$-value indicates the miRNA was assigned a $q$-value of less than 0.35 by the corresponding method.
+The bottom rows report, for each method, the number of selected features with over 1000 citations, under 1000 citations, and the total number selected at the maximum target FDR of 0.35.
+Gene Citations IPSSGB IPSSRF IPSSL1 KOGLM KORF KOL1 DeepPINK SSBoost
+CD38 109641 0.13 0.14
+AKT2 97889 0.13 0.30
+ERBB4 61018 0.45
+CCR3 25401 0.23 0.35
+CD1C 24420 0.38
+WTAP 22418 0.24
+SHMT2 19844 0.30
+AAAS 16226 0.30 0.45
+PAK4 14396 0.38
+SLAMF7 12027 0.24 0.14
+$\geq$ 200 15 14 8 13 2 7 0 0
+$<$ 200 4 1 3 4 1 0 0 0
+Total 19 15 11 17 3 7 0 0
+Table S8.
+RNA-seq and prognosis (ovarian cancer).
+Genes are ordered by citation count.
+A missing $q$-value indicates the gene was assigned a $q$-value of less than 0.5 by the corresponding method.
+The bottom rows report, for each method, the number of selected features with over 200 citations, under 200 citations, and the total number selected at the maximum target FDR of 0.5.
+miRNA Citations IPSSGB IPSSRF IPSSL1 KOGLM KORF KOL1 DeepPINK SSBoost
+miR-155 94172 0.05
+miR-10b 42000 0.21 0.05 0.03 0.35 0.39
+miR-148a 32559 0.13
+miR-335 22423 0.08 0.05 0.50
+miR-15b 21960 0.14 0.05 0.03 0.20 0.38
+miR-424 21680 0.37
+miR-10a 20065 0.21
+miR-224 18932 0.16 0.38
+miR-503 11894 0.14 0.06 0.38
+let-7e 11348 0.25 0.06 0.38
+$\geq$ 100 15 21 6 0 0 14 0 14
+$<$ 100 5 2 3 0 0 10 0 4
+Total 20 23 9 0 0 24 0 18
+Table S9.
+MicroRNAs and prognosis (glioma).
+MiRNAs are ordered by citation count.
+A missing $q$-value indicates the miRNA was assigned a $q$-value of less than 0.5 by the corresponding method.
+The bottom rows report, for each method, the number of selected features with over 100 citations, under 100 citations, and the total number selected at the maximum target FDR of 0.5.
+Gene Citations IPSSGB IPSSRF IPSSL1 KOGLM KORF KOL1 DeepPINK SSBoost
+FOXM1 62538 0.25
+WEE1 26648 0.10 0.06 0.44
+IGFBP2 24482 0.08
+CX3CL1 23044 0.34
+TIMP1 22632 0.44
+SKI 19220 0.12 0.23 0.03 0.44 0.34
+CCNB1 14856 0.10
+CDK9 14558 0.25
+TOP2A 13820 0.03
+PDPN 12929 0.23
+MSN 11866 0.10
+ATF2 11040 0.10
+$\geq$ 500 19 18 12 0 22 24 0 0
+$<$ 500 9 6 7 0 15 36 0 0
+Total 28 24 19 0 37 60 0 0
+Table S10.
+RNA-seq and prognosis (glioma).
+Genes are ordered by citation count.
+A missing $q$-value indicates the gene was assigned a $q$-value of less than 0.5 by the corresponding method.
+The bottom rows report, for each method, the number of selected features with over 500 citations, under 500 citations, and the total number selected at the maximum target FDR of 0.5.
+S4.4  Cross-validation.
+As noted in the main text, we also measure feature selection performance by implementing a 20-fold cross-validation (CV) procedure, described as follows.
+In each of the 20 CV steps, one group of patients is set aside (the test set), and a set of features is selected by each method using the data in the remaining groups (the training set).
+Next, for each method, we construct three predictive models—a linear model, a random forest model, and a gradient boosting model—using only the features selected by that method on the training data.
+Each model is then used to predict responses from the test set, and the smallest of the three prediction errors is recorded (we use mean squared error for regression and $1-\mathrm{accuracy}$ for classification).
+All three models are implemented to ensure that no method has an inherent advantage over another.
+For example, the features selected by IPSSL1 may be better suited to minimizing error in a linear model than those selected by IPSSGB, while those selected by IPSSGB may be better suited to minimizing error in a gradient boosting model than the ones selected by IPSSL1.
+The linear and random forest predictive models are implemented with scikit-learn (Pedregosa et al., 2011) and gradient boosting with XGBoost (Chen and Guestrin, 2016), always with default parameters.
+For continuous responses, in each CV step we subtract the mean of the training responses from all responses, training and test, and scale all responses by the empirical standard deviation of the training responses.
+CV study results are shown in Figures S6S11.
+Each plot contains two subplots.
+The left subplots show the prediction error associated to the features selected by each method at the given target FDR.
+In all studies, we find that the three IPSS methods select features at much lower target FDRs than all of the model-X knockoffs methods.
+The right subplots show the prediction error as a function of the number of features selected by each method.
+Curves for each method in these plots are obtained by varying the target FDR between 0 and 0.5.
+Boruta does not have FDR control parameters and is therefore represented by a single point in these plots.
+In each plot, the dashed black line shows the average error when using all features in the dataset to predict the response variable.
+DeepPINK rarely selects any features and is therefore represented by a single point for better visibility.
+Figure S6. MiRNAs and tumor purity (ovarian cancer).
+Figure S7. MiRNAs and miR-150 (ovarian cancer).
+Figure S8. Genes and AKT2 (ovarian cancer).
+Figure S9. MiRNAs and miR-155 (glioma).
+Figure S10. Genes and prognosis (glioma).
+Figure S11. Genes and FOXM1 (glioma).
+S5.  Sensitivity analyses
+Figures S12S19 show the sensitivity of IPSSGB to the IPSS parameters discussed in Sections 2 and S1.3.
+Data are simulated according to the ovarian cancer RNA-seq simulation design described in Sections 3.3 and S3.2 for both regression and classification.
+The results for IPSSRF were similar and are therefore omitted.
+Figure S12. Different choices of $C$ (regression).
+Figure S13. Different choices of $C$ (classification).
+Figure S14. Different choices of $\delta$ (regression).
+Figure S15. Different choices of $\delta$ (classification).
+Figure S16. Different choices of $B$ (regression).
+Figure S17. Different choices of $B$ (classification).
+Figure S18. Different choices of function (regression).
+Figure S19. Different choices of function (classification).